### v1

# Asymptotics of the maximum likelihood estimator of the location parameter of Pearson Type VII distribution

Abstract.
We study the maximum likelihood estimator of the location parameter of the Pearson Type VII distribution with known scale. We rigorously establish precise asymptotic properties such as strong consistency, asymptotic normality, Bahadur efficiency and asymptotic variance of the maximum likelihood estimator. Our focus is the heavy-tailed case, including the Cauchy distribution. The main difficulty lies in the fact that the likelihood equation may have multiple roots; nevertheless, the maximum likelihood estimator performs well for large samples.
Key words and phrases:
Pearson Type VII distribution, Cauchy distribution, maximum likelihood estimator, strong consistency, asymptotic efficiency, Bahadur efficiency
2020 Mathematics Subject Classification:
Primary 62F12
1. Introduction
The family of Pearson Type VII distributions provides flexible heavy-tailed models. The estimation of its parameters dates back at least to Fisher [7], nearly a century ago, and many researchers have studied it since then; see Johnson, Kotz, and Balakrishnan [9, Section 28] for a thorough survey of results prior to 1994. This class is also known as the location–scale family of Student’s $t$ distributions or of $q$-Gaussian distributions. For estimating the location, the median is a robust alternative to the arithmetic mean; however, it is not asymptotically efficient.
In general, the maximum likelihood estimator is widely regarded as optimal in large samples under standard regularity. Lange, Little, and Taylor [10] proposed a strategy based on maximum likelihood for a general model with errors following the $t$-distribution and applied it to many problems. Under suitable regularity conditions, properties such as strong consistency, asymptotic efficiency, and Bahadur efficiency have been established by many researchers. For location–scale families, it is natural to consider the estimation of the location with known scale. The standard approach is to solve the likelihood equation explicitly or numerically, which often has a unique root. For the Cauchy distribution with known scale, however, the likelihood equation may have multiple roots (see Reeds [12] for precise analysis), and the same phenomenon occurs for the Pearson Type VII distribution. For this reason, alternative estimators of the Cauchy location parameter have been considered. For example, Cohen Freue [6] considered the Pitman estimator for small samples, and Zhang [17] considered an empirical Bayes estimator. Nevertheless, this does not represent a failure of the maximum likelihood estimator itself. Indeed, Bai and Fu [2] established its Bahadur efficiency.
In this paper, we deal with not only the Cauchy distribution but also the Pearson Type VII distribution and our focus is the maximum likelihood estimator. Some references on the maximum likelihood estimator of the Pearson Type VII distribution are Borwein and Gabor [5], Tiku and Suresh [15], and Vaughan [16]. We provide mathematically rigorous proofs of strong consistency, asymptotic efficiency, and Bahadur efficiency for the maximum likelihood estimator. Our approach does not analyze the likelihood equation directly. We show that the asymptotic properties of the maximum likelihood estimator mirror those for the arithmetic mean of independent and identically distributed (i.i.d.) random variables with finite variance. Asymptotically, the maximum likelihood estimator for the Pearson Type VII distribution behaves well.
Now we state the framework and the main result. Let $m>1/2$, which covers the heavy-tailed regime of primary interest. Let $\textup{PVII}_{m}(\mu,\sigma)$ be the Pearson Type VII distribution with location $\mu\in\mathbb{R}$ and scale $\sigma>0$. Then, the probability density function of $\textup{PVII}_{m}(\mu,\sigma)$ is given by
$f(x)=c_{m}\frac{1}{\sigma}\left(1+\left(\frac{x-\mu}{\sigma}\right)^{2}\right)^{-m},$
where $c_{m}$ is the normalizing constant, specifically, $c_{m}:=\left(\int_{\mathbb{R}}(1+x^{2})^{-m}dx\right)^{-1}$. The case that $m=1$ is the Cauchy distribution.
We consider the maximum likelihood estimator of the location parameter of the Pearson Type VII distribution with known scale. We can assume that $\sigma=1$. Let $(X_{n})_{n\geq 1}$ be i.i.d. random variables on a probability space $(\Omega,\mathcal{F},P)$ following $\textup{PVII}_{m}(\theta,1)$. Let $\hat{\theta}_{n}$ be the maximum likelihood estimator of the location parameter from a sample $(X_{1},\dots,X_{n})$ of size $n$. Let $\hat{\theta}_{n}(x_{1},\dots,x_{n})$ be a measurable function on $\mathbb{R}^{n}$ which maximizes the function $\theta\mapsto\prod_{i=1}^{n}f(x_{i}-\theta)$. Such a function exists by virtue of the measurable selection theorem. Then, let $\hat{\theta}_{n}:=\hat{\theta}_{n}(X_{1},\dots,X_{n})$.
Our first main result is strong consistency.
Theorem 1.1 (Strong consistency).
$\lim_{n\to\infty}\hat{\theta}_{n}=\theta,\textup{ $P$-a.s.}$
We show this using the concept of the Fréchet mean.
Once we see the strong consistency, it is natural to consider the asymptotic normality. We denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$ by $N(\mu,\sigma^{2})$.
Theorem 1.2 (Asymptotic normality).
$\left(\sqrt{n}(\hat{\theta}_{n}-\theta)\right)_{n}$ converges to $N\left(0,\dfrac{m+1}{m(2m-1)}\right)$ in distribution as $n\to\infty$.
By Remark 3.5 below, $I(\theta)=\dfrac{m(2m-1)}{m+1}$, where $I(\theta)$ is the Fisher information for a single observation.
We can also see the following behavior, although it is not often considered in mathematical statistics.
Theorem 1.3 (Law of the iterated logarithm).
$\limsup_{n\to\infty}\sqrt{\frac{n}{\log\log n}}(\hat{\theta}_{n}-\theta)=\sqrt{\frac{2(m+1)}{m(2m-1)}},\textup{ $P$-a.s.}$
For the proof, we use the technique of the deviation mean of i.i.d. random variables investigated by Barczy and Páles [3] with some modifications.
The following extends the result of Bai and Fu [2], who considered the Cauchy distribution, to the Pearson Type VII distribution.
Theorem 1.4 (Bahadur efficiency and moderate deviation).
$\limsup_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\limsup_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\leq-\frac{m(2m-1)}{2(m+1)}.$ (1.1)
$\liminf_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\liminf_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\geq-\frac{m(2m-1)}{2(m+1)}.$ (1.2)
(ii) For every sequence $(\lambda_{n})_{n}$ of positive numbers satisfying $\lim_{n\to\infty}\lambda_{n}=\infty$ and $\lim_{n\to\infty}\lambda_{n}/n^{1/2}=0$ and every $\epsilon>0$,
$\lim_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon/\lambda_{n}\right)}{n/\lambda_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
This assertion implies Theorem 1.1 and its proof does not depend on Theorem 1.1. However, we can show Theorem 1.1 much more easily than the proof of this assertion. For the proof, we follow the strategy of [2].
It is worth investigating the probability that the estimator deviates significantly from the true value. Hereafter, $\mathbb{N}=\{1,2,\dots\}$.
Theorem 1.5 (Integrability).
There exist positive constants $c_{m},r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
$P\left(\left|\hat{\theta}_{n}-\theta\right|>r\right)\leq r^{-c_{m}n}.$
In particular, $\hat{\theta}_{n}\in L^{c_{m}n-1}(\Omega,\mathcal{F},P)$ for $n\geq N_{m}$.
We show this by modifying several estimates in the proof of Theorem 1.4.
The Cramér-Rao inequality states that for each $n\geq 1$,
$nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]\geq\frac{1}{I(\theta)}.$
By this and Theorem 1.2, it is natural to consider the large-sample asymptotics of $nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$.
Theorem 1.6 (Variance asymptotics).
$\lim_{n\to\infty}nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]=\frac{m+1}{m(2m-1)}.$
This is consistent with [9, (28.61c)]. We give a mathematically rigorous proof of it. The proof is technically involved and we use Theorems 1.2 and 1.5.
In the following sections, we present proofs of these assertions. In the final section, we give numerical computations of $nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$.
In the proofs of these results, we can assume that $\theta=0$ without loss of generality. The parameter $m$ remains fixed throughout. Many constants will appear. When a constant depends only on $m$, we indicate this by attaching $m$ to its index; otherwise we omit it even if it depends on $m$.
2. Proof of Theorem 1.1
We prove Theorem 1.1 by following the strategy of [4, Section 3.2]. One of our goals is to establish [4, Theorem 3.3] in the case where the loss function is replaced [^1] In [4, Theorem 3.3], the loss function is given by the map $u\mapsto u^{\alpha}$ for $\alpha\in(0,1)$. with the map $u\mapsto\log(1+u^{2})$.
Let
$L_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}\log(1+(X_{i}-t)^{2}),\ t\in\mathbb{R}.$
Let $\nu_{m}$ be the Borel probability measure of the Pearson Type VII distribution $\textup{PVII}_{m}(0,1)$, that is,
$\nu_{m}(dx)=c_{m}\left(1+x^{2}\right)^{-m}dx.$
Lemma 2.1.
There exists a positive constant $c_{m,1}$ depending only on $m$ such that $P$-a.s. $\omega$, there exists $N_{1}(\omega)\in\mathbb{N}$ such that for every $n>N_{1}(\omega)$ and every $t\in\mathbb{R}$ with $|t|\geq 2$,
$L_{n}(t)(\omega)\geq\frac{c_{m,1}}{4}(\log(1+t^{2})-2\log 2).$
Proof.
Applying the inequality
$\log(1+x^{2})+\log(1+y^{2})\geq\frac{1}{2}\log(1+(x+y)^{2}),\ x,y\in\mathbb{R},$
to $(x,y)=(X_{i}(\omega)-t,X_{i}(\omega))$, we see that
$L_{n}(t)(\omega)\geq\frac{1}{n}\sum_{i=1}^{n}\log(1+(X_{i}(\omega)-t)^{2}){\bf 1}_{[-1,1]}(X_{i}(\omega))$
$\geq\left(\frac{1}{2}\log(1+t^{2})-\log 2\right)\frac{1}{n}\sum_{i=1}^{n}{\bf 1}_{[-1,1]}(X_{i}(\omega)).$
By the strong law of large numbers,
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}{\bf 1}_{[-1,1]}(X_{i}(\omega))=\nu_{m}([-1,1])>0,\ \textup{ $P$-a.s.$\omega$}.$
We have the assertion for $c_{m,1}:=\nu_{m}([-1,1])$. ∎
Denote the empirical distribution of $(X_{i}(\omega))_{i=1}^{n}$ by $Q_{n}(\omega)$, specifically,
$Q_{n}(\omega)=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}(w)}.$
Lemma 2.2 (boundedness of minimizers).
There exists a positive constant $r_{m,1}$ depending only on $m$ such that $P$-a.s. $\omega$, there exists $N_{2}(\omega)\in\mathbb{N}$ such that for every $n>N_{2}(\omega)$, $C_{Q_{n}(\omega)}\subset[-r_{m,1},r_{m,1}]$.
Proof.
By Lemma 2.1, there exists an event $\Omega_{1}$ such that $P(\Omega_{1})=1$ and for every $\omega\in\Omega_{1}$, there exists $N_{1}(\omega)\in\mathbb{N}$ such that for every $n>N_{1}(\omega)$ and every $t\in\mathbb{R}$ with $|t|\geq 2$,
$L_{n}(t)(\omega)\geq\frac{c_{m,1}}{4}(\log(1+t^{2})-2\log 2).$
Assume that $\omega\in\Omega_{1}$, $t_{n}\in C_{Q_{n}(\omega)}$ and $|t_{n}|\geq 2$. Then, for every $n>N_{1}(\omega)$,
$L_{n}(0)(\omega)\geq L_{n}(t_{n})(\omega)\geq\frac{c_{m,1}}{4}(\log(1+t_{n}^{2})-2\log 2)$
By the strong law of large numbers, there exists an event $\Omega_{2}\subset\Omega_{1}$ such that $P(\Omega_{2})=1$ and for every $\omega\in\Omega_{2}$,
$\lim_{n\to\infty}L_{n}(0)(\omega)=F_{m}(0)<+\infty.$
In particular, there exists $N_{2}(\omega)>N_{1}(\omega)$ such that for every $n>N_{2}(\omega)$,
$L_{n}(0)(\omega)\leq 1+F_{m}(0).$
Hence, there exists a constant $r_{m,1}>1$ such that for every $\omega\in\Omega_{2}$ and $n>N_{2}(\omega)$, $|t_{n}|<r_{m,1}$. ∎
For a Borel probability measure $\nu$ on $\mathbb{R}$, let the expected loss function be
$F_{\nu}(t):=\int_{\mathbb{R}}\log(1+(x-t)^{2})d\nu(x),\ t\in\mathbb{R},$
and the mean set be
$C_{\nu}:=\left\{t\in\mathbb{R}\middle|\min_{s\in\mathbb{R}}F_{\nu}(s)=F_{\nu}(t)\right\}.$
For $\nu=Q_{n}(\omega)$, $C_{Q_{n}(\omega)}$ is called the Fréchet mean set. It holds that $\hat{\theta}_{n}(\omega)\in C_{Q_{n}(\omega)}$.
Another goal of this section is to show that the mean set $C_{\nu_{m}}$ is a singleton, which will be established in Lemma 2.5 below.
For notational convenience,
$F_{m}(t):=F_{\nu_{m}}(t)=\int_{\mathbb{R}}\log(1+(x-t)^{2})\nu_{m}(dx).$ (2.1)
This is the expected loss function.
Lemma 2.3 (a.s. pointwise convergence).
$P$-a.s., it holds that every $t\in\mathbb{R}$,
$\lim_{n\to\infty}L_{n}(t)=F_{m}(t).$ (2.2)
Proof.
By the strong law of large numbers, for every fixed $t\in\mathbb{R}$, (2.2) holds a.s. Now, use the rational approximation and use the Lipschitz continuity [^2] This estimate works well if $x$ and $y$ are close. If $x$ or $y$ is large, the bound can be very loose. of $\log(1+x^{2})$, specifically,
$\left|\log(1+x^{2})-\log(1+y^{2})\right|\leq\left||x|-|y|\right|\leq|x-y|.$ (2.3)
The following is the uniform law of large numbers.
Lemma 2.4.
$P$-a.s., it holds that for every compact subset $K$ of $\mathbb{R}$,
$\lim_{n\to\infty}\max_{t\in K}\left|L_{n}(t)-F_{m}(t)\right|=0.$
Proof.
By Lemma 2.1, there exists an event $\Omega_{3}$ such that $P(\Omega_{3})=1$ and for every $\omega\in\Omega_{3}$, (2.2) holds for every $t\in\mathbb{R}$.
Let $\omega\in\Omega_{3}$. Let $\epsilon>0$ arbitrarily. Then, by (2.3), for each $t_{1},t_{2}\in\mathbb{R}$ with $|t_{1}-t_{2}|<\epsilon/4$,
$\left|F_{m}(t_{1})-F_{m}(t_{2})\right|\leq|t_{1}-t_{2}|\leq\frac{\epsilon}{4}.$
Let $u_{1},\cdots,u_{\ell}$ be points in $K$ such that $K\subset\cup_{j=1}^{\ell}(u_{j}-\epsilon/4,u_{j}+\epsilon/4)$. Then, by Lemma 2.3, there exists $N_{3}(\omega)\in\mathbb{N}$ such that for every $n>N_{3}(\omega)$,
$\max_{1\leq j\leq\ell}\left|L_{n}(u_{j})(\omega)-F_{m}(u_{j})\right|<\frac{\epsilon}{4}.$
Then, by (2.3), if $t\in K$ and $|t-u_{j}|<\epsilon/4$, then, for every $n>N_{3}(\omega)$,
$\left|L_{n}(t)(\omega)-F_{m}(t)\right|\leq\frac{\epsilon}{2}+\left|L_{n}(u_{j})(\omega)-L_{n}(t)(\omega)\right|\leq\epsilon.$
Lemma 2.5.
$F_{m}$ is strictly decreasing on $(-\infty,0)$ and strictly increasing on $(0,\infty)$. In particular, $C_{\nu_{m}}=\{0\}$.
Proof.
By the Lebesgue convergence theorem, we see that
$F_{m}^{\prime}(t)=-2c_{m}\int_{\mathbb{R}}\frac{x-t}{(1+(x-t)^{2})(1+x^{2})^{m}}dx.$ (2.4)
By change of variables,
$\int_{\mathbb{R}}\frac{x-t}{(1+(x-t)^{2})(1+x^{2})^{m}}dx=\int_{\mathbb{R}}\frac{x}{(1+x^{2})(1+(x+t)^{2})^{m}}dx$
$\int_{0}^{\infty}\frac{x}{(1+x^{2})(1+(x+t)^{2})^{m}}dx+\int_{-\infty}^{0}\frac{x}{(1+x^{2})(1+(x+t)^{2})^{m}}dx$
$\int_{0}^{\infty}\frac{x}{1+x^{2}}\left(\frac{1}{(1+(x+t)^{2})^{m}}-\frac{1}{(1+(x-t)^{2})^{m}}\right)dx.$
The last integral is positive if $t<0$, is zero if $t=0$, and is negative if $t>0$. Hence, the sign of $F_{m}^{\prime}(t)$ is equal to the sign of $t$, and hence, $F_{m}(t)$ takes its minimum only at $t=0$. ∎
For a non-empty subset $A$ of $\mathbb{R}$, let
$d(x,A):=\inf\left\{|x-y|:y\in A\right\}.$
Lemma 2.6.
Let $\varphi$ be a continuous function on $\mathbb{R}$ such that $\lim_{|z|\to\infty}\varphi(z)=\infty$. Let $C_{\varphi}:=\left\{x\in\mathbb{R}:\min_{t\in\mathbb{R}}\varphi(t)=\varphi(x)\right\}$. Then, for every $\epsilon>0$, there exists $\delta>0$ such that for every $x\in\mathbb{R}$ with $\varphi(x)\leq\min_{t\in\mathbb{R}}\varphi(t)+\delta$, $d(x,C_{\varphi})<\epsilon$.
Proof.
We show this by contradiction. Assume that there exists $\epsilon_{0}>0$ such that for every $n\in\mathbb{N}$, there exists $x_{n}\in\mathbb{R}$ such that $\varphi(x_{n})\leq\min_{t\in\mathbb{R}}\varphi(t)+1/n$ and $d(x_{n},C_{\varphi})\geq\epsilon_{0}$. Since $\sup_{n\in\mathbb{N}}\varphi(x_{n})<+\infty$, by the assumption of $\varphi$, $(x_{n})_{n}$ is a bounded sequence. Then, there exist a subsequence $(x_{n_{k}})_{k}$ and a point $z\in\mathbb{R}$ such that $x_{n_{k}}\to z,k\to\infty$. By the continuity of $\varphi$, $\varphi(z)=\lim_{k\to\infty}\varphi(x_{n_{k}})=\min_{t\in\mathbb{R}}\varphi(t)$. Hence, $z\in C_{\varphi}$. Now it suffices to recall that $|x_{n_{k}}-z|\geq d(x_{n_{k}},C_{\varphi})\geq\epsilon_{0}$ for each $k$. ∎
Proposition 2.7 (confinement of minimizers).
$P$-a.s. $\omega$, it holds that for every $\epsilon>0$, there exists $N_{4}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n>N_{4}(\omega,\epsilon)$, $C_{Q_{n}(\omega)}\subset[-\epsilon,\epsilon]$.
Proof.
By applying Lemma 2.2 and Lemma 2.4 to $K=[-r_{m,1},r_{m,1}]$, it holds that for every $\omega\in\Omega_{3}$ and $t_{n}\in C_{Q_{n}(\omega)}$,
$\lim_{n\to\infty}\left|L_{n}(t_{n})(\omega)-F_{m}(t_{n})\right|=0.$
Let $\epsilon>0$. Then, there exists $N_{4}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n>N_{4}(\omega,\epsilon)$ and $t_{n}\in C_{Q_{n}(\omega)}$,
$F_{m}(t_{n})\leq L_{n}(t_{n})(\omega)+\frac{\epsilon}{4}$
and
$F_{m}(0)\geq L_{n}(0)(\omega)-\frac{\epsilon}{4},$
in particular,
$L_{n}(t_{n})\leq F_{m}(0)+\epsilon.$
Now apply Lemma 2.6 to $\varphi=F_{m}$ and Lemma 2.5. ∎
By Proposition 2.7, we obtain Theorem 1.1.
Remark 2.8.
Recently, Schötz [13] gave precise analysis for the Fréchet mean. His approach uses a general ergodic theorem and differs from the above approach.
3. Proof of Theorem 1.2
We follow the strategy of [3, Section 4]. Let
$D(x,t):=\frac{x-t}{1+(x-t)^{2}},\ x,t\in\mathbb{R}.$
Let
$D_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}D(X_{i},t).$
Then, $2D_{n}(t)\equiv S_{n}^{\prime}(t)$ and hence, the likelihood equation is $D_{n}(t)=0$.
We first show that $\hat{\theta}_{n}>t$ holds if and only if $D_{n}(t)>0$ on an event with high probability. Since $D(x,t)$ is not monotone with respect to $t$ on $\mathbb{R}$, we cannot apply the result of [3] directly and need to “localize” it. The arguments in Section 2 are not sufficient for this goal, since the possibility that $|C_{Q_{n}}\cap[-T,T]|\geq 2$ has not yet been excluded.
Let
$G_{m}(t):=E\left[D(X_{1},t)\right]=\int_{\mathbb{R}}D(x,t)\nu_{m}(dx).$
Then, by (2.4), $2G_{m}(t)\equiv F_{m}^{\prime}(t)$ and hence, $G_{m}(-t)>0>G_{m}(t)$ for every $t>0$.
We show that
Lemma 3.1.
For every $\epsilon>0$, there exists a positive constant $c_{\epsilon,1}$ depending on $\epsilon$ such that for every $n\geq 1$,
$P\left(\max_{t\in[-1,1]}\left|D_{n}(t)-G_{m}(t)\right|>\epsilon\right)\leq c_{\epsilon,1}\exp\left(-\frac{\epsilon^{2}}{2}n\right).$
In particular,
$\lim_{n\to\infty}\max_{t\in[-1,1]}\left|D_{n}(t)-G_{m}(t)\right|=0,\ \textup{ $P$-a.s.}$
The constant $c_{\epsilon,1}$ is independent of $m$.
Proof.
Let $t\in[-1,1]$. Let $Y_{i}:=D(X_{i},t)-G_{m}(t)$. Since $|D(X_{i},t)|\leq 1/2$ and $|G_{m}(t)|\leq 1/2$, $(Y_{i})_{i}$ are i.i.d., $|Y_{i}|\leq 1$ and $E[Y_{i}]=0$. By the Azuma-Hoeffding inequality (see Petrov [11, 2.6.2]),
$P\left(\left|D_{n}(t)-G_{m}(t)\right|>\epsilon\right)=P\left(\left|\sum_{i=1}^{n}Y_{i}\right|>n\epsilon\right)\leq 2\exp\left(-\frac{\epsilon^{2}}{2}n\right).$
Let $\mathcal{D}_{N}:=\{\ell/N:-N\leq\ell\leq N\}$ for $N\in\mathbb{N}$. Since $t\mapsto D(x,t)$ is Lipschitz continuous with the Lipschitz constant $1$,
$\max_{t\in[-1,1]}\left|D_{n}(t)-G_{m}(t)\right|\leq\max_{t\in\mathcal{D}_{N}}\left|D_{n}(t)-G_{m}(t)\right|+\frac{2}{N}.$
Hence, for $N>4/\epsilon$ and $n\geq 1$,
$P\left(\max_{t\in[-1,1]}\left|D_{n}(t)-G_{m}(t)\right|>\epsilon\right)\leq P\left(\max_{t\in\mathcal{D}_{N}}\left|D_{n}(t)-G_{m}(t)\right|>\frac{\epsilon}{2}\right)$
$\leq\sum_{t\in\mathcal{D}_{N}}P\left(\left|D_{n}(t)-G_{m}(t)\right|>\frac{\epsilon}{2}\right)\leq 2(2N+1)\exp\left(-\frac{\epsilon^{2}}{2}n\right).$
Now use the Borel-Cantelli lemma and then let $\epsilon\to+0$, and we obtain the a.s. convergence. ∎
We see that
$\partial_{t}D(x,t)=\frac{(x-t)^{2}-1}{(1+(x-t)^{2})^{2}}.$
Lemma 3.2.
There exists a constant $r_{m,2}\in(0,1)$ such that $G_{m}^{\prime}(t)<0$ for every $t\in[-r_{m,2},r_{m,2}]$.
Proof.
By the Lebesgue convergence theorem, we see that
$G_{m}^{\prime}(t)=\int_{\mathbb{R}}\partial_{t}D(x,t)\nu_{m}(dx).$
By the Lebesgue convergence theorem, $G_{m}^{\prime}$ is continuous. Hence, it suffices to show that $G_{m}^{\prime}(0)<0$.
By the change of variables $x=\tan\theta$,
$\int_{\mathbb{R}}\frac{x^{2}-1}{(1+x^{2})^{2+m}}dx=-\int_{-\pi/2}^{\pi/2}\cos^{2m}\theta\cos(2\theta)d\theta=-2\int_{0}^{\pi/2}\cos^{2m}\theta\cos(2\theta)d\theta.$
We see that
$\int_{0}^{\pi/2}\cos^{2m}\theta\cos(2\theta)d\theta=\int_{0}^{\pi/4}\cos^{2m}\theta\cos(2\theta)d\theta+\int_{\pi/4}^{\pi/2}\cos^{2m}\theta\cos(2\theta)d\theta$
$\int_{0}^{\pi/4}\cos^{2m}\theta\cos(2\theta)d\theta-\int_{0}^{\pi/4}\cos^{2m}\left(\frac{\pi}{2}-\theta\right)\cos(2\theta)d\theta>0.$
We also deal with the derivatives of $D_{n}(t)$ and $G_{m}(t)$ with respect to $t$. The following corresponds to [2, (3.32)].
Lemma 3.3.
For every $\epsilon>0$, there exists a positive constant $c_{\epsilon,2}$ depending on $\epsilon$ such that for every $n\geq 1$,
$P\left(\max_{t\in[-1,1]}\left|D_{n}^{\prime}(t)-G_{m}^{\prime}(t)\right|>\epsilon\right)\leq c_{\epsilon,2}\exp\left(-\frac{\epsilon^{2}}{12}n\right).$
In particular,
$\lim_{n\to\infty}\max_{t\in[-r_{m,2},r_{m,2}]}\left|D_{n}^{\prime}(t)-G_{m}^{\prime}(t)\right|=0,\ \textup{ $P$-a.s.}$
As in Lemma 3.3, the constant $c_{\epsilon,2}$ is also independent of $m$.
Proof.
By (3.7) below, $|\partial_{t}^{2}D(x,t)|\leq 3$, and hence, the map $t\mapsto\partial_{t}D(x,t)$ is Lipschitz continuous with the Lipschitz constant $3$. Let $Y^{\prime}_{i}:=\partial_{t}D(X_{i},t)-G_{m}^{\prime}(t)$. Since $|\partial_{t}D(X_{i},t)|\leq 1$ and $|G_{m}^{\prime}(t)|\leq 1$, $\left(Y^{\prime}_{i}\right)_{i}$ are i.i.d., $|Y_{i}^{\prime}|\leq 2$ and $E[Y_{i}^{\prime}]=0$. Therefore, we can show this assertion as in the proof of Lemma 3.1. ∎
We remark that $C_{Q_{n}(\omega)}\neq\emptyset$ and
$C_{Q_{n}(\omega)}\subset\left\{t\in\mathbb{R}\middle|D_{n}(t)(\omega)=0\right\}.$
Proposition 3.4.
$P$-a.s. $\omega$, there exists $N_{5}(\omega)\in\mathbb{N}$ such that for every $n>N_{5}(\omega)$, $|C_{Q_{n}(\omega)}\cap[-r_{m,2},r_{m,2}]|=1$.
Proof.
By Proposition 2.7, it holds that $P$-a.s. $\omega$, for $n\geq N_{4}(\omega,r_{m,2})$, $|C_{Q_{n}(\omega)}\cap[-r_{m,2},r_{m,2}]|=|C_{Q_{n}(\omega)}|\geq 1$.
Let
$c_{m,2}:=\frac{1}{2}\min_{t\in[-r_{m,2},r_{m,2}]}-G_{m}^{\prime}(t),$
which is positive by Lemma 3.2.
By Lemma 3.3, it holds that $P$-a.s. $\omega$, there exists $N_{6}(\omega)\in\mathbb{N}$ such that for every $n>N_{6}(\omega)$,
$\max_{t\in[-r_{m,2},r_{m,2}]}D_{n}^{\prime}(t)(\omega)\leq-c_{m,2},$
in particular, $D_{n}(t)(\omega)$ is strictly decreasing in $t$ on $r_{m,2},r_{m,2}]$.
Furthermore, by Lemma 3.1, it holds that $P$-a.s. $\omega$, there exists $N_{7}(\omega)\in\mathbb{N}$ such that for every $n>N_{7}(\omega)$,
$D_{n}(-r_{m,2})(\omega)>0>D_{n}(r_{m,2})(\omega).$
By the intermediate value theorem, it holds that $P$-a.s. $\omega$, there exists $N_{8}(\omega)\in\mathbb{N}$ such that for every $n>N_{8}(\omega)$,
$\left|\left\{t\in[-r_{m,2},r_{m,2}]\middle|D_{n}(t)(\omega)=0\right\}\right|=1,$
which implies $|C_{Q_{n}(\omega)}\cap[-r_{m,2},r_{m,2}]|\leq 1$. ∎
Let $\mathcal{A}_{n,1}$ be the event that $D_{n}(-r_{m,2})>0>D_{n}(r_{m,2})$. Let $\mathcal{A}_{n,2}$ be the event that $D_{n}^{\prime}(t)\leq-c_{m,2}/2$ for every $t\in[-r_{m,2},r_{m,2}]$. Let $\mathcal{A}_{n,3}$ be the event that $|C_{Q_{n}}|=1$ and $\hat{\theta}_{n}\in[-r_{m,2}/2,r_{m,2}/2]$. Let
$\mathcal{A}_{n}:=\mathcal{A}_{n,1}\cap\mathcal{A}_{n,2}\cap\mathcal{A}_{n,3}.$
Let $\widetilde{\mathcal{A}_{i}}:=\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n,i},i=1,2,3$.
By Lemma 3.1, $P\left(\widetilde{\mathcal{A}_{1}}\right)=1$. By Lemma 3.3, $P\left(\widetilde{\mathcal{A}_{2}}\right)=1$. By Propositions 2.7 and 3.4, $P\left(\widetilde{\mathcal{A}_{3}}\right)=1$. Since $\widetilde{\mathcal{A}_{1}}\cap\widetilde{\mathcal{A}_{2}}\cap\widetilde{\mathcal{A}_{3}}=\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}$, $P\left(\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}\right)=1$, and in particular, $\lim_{n\to\infty}P(\mathcal{A}_{n})=1$.
For every $t\in(-r_{m,2}/2,r_{m,2}/2)$, on $\mathcal{A}_{n}$, $\hat{\theta}_{n}<t$ if and only if $D_{n}(t)<0$.
Let $y\in\mathbb{R}$. Then,
$\lim_{n\to\infty}P\left(\sqrt{n}\hat{\theta}_{n}<y\right)-P\left(\left\{\sqrt{n}\hat{\theta}_{n}<y\right\}\cap\mathcal{A}_{n}\right)=0,$
and
$\lim_{n\to\infty}P\left(D_{n}\left(\frac{y}{\sqrt{n}}\right)<0\right)-P\left(\left\{D_{n}\left(\frac{y}{\sqrt{n}}\right)<0\right\}\cap\mathcal{A}_{n}\right)=0.$
Since
$P\left(\left\{\sqrt{n}\hat{\theta}_{n}<y\right\}\cap\mathcal{A}_{n}\right)=P\left(\left\{D_{n}\left(\frac{y}{\sqrt{n}}\right)<0\right\}\cap\mathcal{A}_{n}\right)$
for every $n$ satisfying that $n>4y^{2}$,
$\lim_{n\to\infty}P\left(\sqrt{n}\hat{\theta}_{n}<y\right)-P\left(D_{n}\left(\frac{y}{\sqrt{n}}\right)<0\right)=0.$
Hence, it suffices to show that
$\lim_{n\to\infty}P\left(D_{n}\left(\frac{y}{\sqrt{n}}\right)<0\right)=\int_{-\infty}^{y}\varphi_{m}(t)dt,$ (3.1)
where $\varphi_{m}$ is the density function of the distribution $N\left(0,\frac{m+1}{m(2m-1)}\right)$.
It holds that
$\sqrt{n}D_{n}\left(\frac{y}{\sqrt{n}}\right)=\sqrt{n}D_{n}\left(0\right)+\frac{y}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0)$
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right).$
By the symmetry,
$E[D(X_{1},0)]=c_{m}\int_{\mathbb{R}}\frac{x}{(1+x^{2})^{1+m}}dx=0.$ (3.2)
By the change of variables $x=\tan\theta$,
$E\left[D(X_{1},0)^{2}\right]=c_{m}\int_{\mathbb{R}}\frac{x^{2}}{(1+x^{2})^{2+m}}dx=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}=\frac{2m-1}{4m(m+1)},$ (3.3)
where $B(\cdot,\cdot)$ is the beta function. Hence,
$\sqrt{n}D_{n}\left(0\right)\Rightarrow N\left(0,\frac{2m-1}{4m(m+1)}\right),\ n\to\infty,$ (3.4)
where $\Rightarrow$ denotes the convergence in distribution.
It holds that
$E\left[|\partial_{t}D(X_{1},0)|\right]\leq c_{m}\int_{\mathbb{R}}\frac{1}{(1+x^{2})^{1+m}}dx<\infty,$
and
$E\left[\partial_{t}D(X_{1},0)\right]$ $=c_{m}\int_{\mathbb{R}}\frac{x^{2}-1}{(1+x^{2})^{2+m}}dx$
$=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}-\frac{B(1/2,m+3/2)}{B(1/2,m-1/2)}=-\frac{2m-1}{2(m+1)}.$ (3.5)
Hence, by the strong law of large numbers,
$\lim_{n\to\infty}\frac{y}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0)=-\frac{2m-1}{2(m+1)}y,\ \textup{ $P$-a.s.}$ (3.6)
Since
$\partial_{t}^{2}D(x,t)=\frac{2(x-t)((x-t)^{2}-3)}{(1+(x-t)^{2})^{3}},$ (3.7)
$\max_{x,t\in\mathbb{R}}\left|\partial_{t}^{2}D(x,t)\right|=\max_{y\in\mathbb{R}}\frac{2|y||y^{2}-3|}{(1+y^{2})^{3}}=:C_{1}<\infty.$
By this and the mean value theorem, it holds [^3] This holds without any exceptional set. that
$\left|D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right|\leq C_{1}\frac{|y|^{2}}{n}.$ (3.8)
Hence,
$\lim_{n\to\infty}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right)=0,\ \textup{ $P$-a.s.}$ (3.9)
By (3.4), (3.6), (3.9) and Slutsky’s theorem,
$\sqrt{n}D_{n}\left(\frac{y}{\sqrt{n}}\right)\Rightarrow N\left(-\frac{2m-1}{2(m+1)}y,\frac{2m-1}{4m(m+1)}\right),\ n\to\infty.$
We see that (3.1) holds.
Remark 3.5.
(i) By (3.3), the Fisher information is given by
$I(0)=E\left[\left(\frac{\partial}{\partial t}\log f(X_{1},t)\Bigg|_{t=0}\right)^{2}\right]=4m^{2}E\left[D(X_{1},0)^{2}\right]=\frac{m(2m-1)}{m+1}.$
(ii) The likelihood equation $D_{n}(t)=0$ does not depend on the parameter $m$. For $m=1$, [12] shows that for each $k\geq 0$,
$\lim_{n\to\infty}P\left(\left|\{t\in\mathbb{R}|D_{n}(t)=0\}\right|=2k+1\right)=\exp\left(-\frac{1}{\pi}\right)\frac{1}{k!\pi^{k}}.$
Here $c_{1}=1/\pi$, and we conjecture that for each $m>1/2$ and $k\geq 0$,
$\lim_{n\to\infty}P\left(\left|\{t\in\mathbb{R}|D_{n}(t)=0\}\right|=2k+1\right)=\exp\left(-c_{m}\right)\frac{c_{m}^{k}}{k!}.$
4. Proof of Theorem 1.3
We follow the strategy of [3, Section 5]. As in the case of Theorem 1.2, we cannot apply the result of [3] directly and need to modify several parts.
Let $\phi(n):=\sqrt{2n\log\log n}$. Let
$R_{n}:=\frac{1}{\phi(n)}\sum_{i=1}^{n}D(X_{i},0),$
$S_{n}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0),$
and
$T_{n}^{(\gamma)}:=\frac{1}{\phi(n)}\sum_{i=1}^{n}\left(D\left(X_{i},\gamma\frac{\phi(n)}{n}\right)-D(x_{i},0)-\gamma\frac{\phi(n)}{n}\partial_{t}D(X_{i},0)\right).$
Then,
$R_{n}+\gamma S_{n}+T_{n}^{(\gamma)}=\frac{1}{\phi(n)}\sum_{i=1}^{n}D\left(X_{i},\gamma\frac{\phi(n)}{n}\right).$ (4.1)
Since $\max_{x,t\in\mathbb{R}}|D(x,t)|\leq 1/2$, by the Kolmogorov law of the iterated logarithm, there exists an event $\Omega_{4}$ such that $P(\Omega_{4})=1$ and for every $\omega\in\Omega_{4}$,
$\limsup_{n\to\infty}R_{n}(\omega)=\sqrt{\frac{2m-1}{4m(m+1)}}.$ (4.2)
By the strong law of large numbers, there exists an event $\Omega_{5}$ such that $P(\Omega_{5})=1$ and for every $\omega\in\Omega_{5}$,
$\lim_{n\to\infty}S_{n}(\omega)=-\frac{2m-1}{2(m+1)}.$ (4.3)
Let $\Omega_{6}:=\Omega_{4}\cap\Omega_{5}\cap\left(\cup_{N\geq 1}\cap_{n\geq N}\mathcal{A}_{n}\right)$. Then, $P\left(\Omega_{6}\right)=1$.
By the uniform estimate (3.8), for every $\omega\in\Omega_{6}$ and every $\gamma\in\mathbb{R}$,
$\lim_{n\to\infty}T_{n}^{(\gamma)}(\omega)=0.$ (4.4)
For notational convenience, let $\sigma_{m}:=\sqrt{\dfrac{m+1}{m(2m-1)}}$. Let $\omega\in\Omega_{6}$ and $\epsilon>0$. Then, there exists $N_{9}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{9}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)<(\sigma_{m}+\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}+\epsilon)\frac{\phi(n)}{n}\right)<0$ holds. By (4.1), this is equivalent to
$R_{n}(\omega)+(\sigma_{m}+\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}+\epsilon)}(\omega)<0.$ (4.5)
By (4.2), (4.3) and (4.4), there exists $N_{10}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{10}(\omega,\epsilon)$, (4.5) holds. Hence, for every $n\geq\max\{N_{9}(\omega,\epsilon),N_{10}(\omega,\epsilon)\}$, $\hat{\theta}_{n}(\omega)<(\sigma_{m}+\epsilon)\dfrac{\phi(n)}{n}$ holds and hence,
$\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\leq\sigma_{m}+\epsilon.$
By letting $\epsilon\to 0$,
$\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\leq\sigma_{m}.$ (4.6)
We can show the lower bound in the same manner. There exists $N_{11}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{11}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)>(\sigma_{m}-\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}-\epsilon)\frac{\phi(n)}{n}\right)>0$ holds. By (4.1), this is equivalent to
$R_{n}(\omega)+(\sigma_{m}-\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}-\epsilon)}(\omega)>0.$ (4.7)
By (4.2), (4.3) and (4.4), (4.7) holds for infinitely many $n$. Hence, $\hat{\theta}_{n}(\omega)>(\sigma_{m}-\epsilon)\dfrac{\phi(n)}{n}$ holds for infinitely many $n$, and hence,
$\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\geq\sigma_{m}-\epsilon.$
By letting $\epsilon\to 0$,
$\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\geq\sigma_{m}.$ (4.8)
By (4.6) and (4.8),
$\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}=\sigma_{m}.$
This completes the proof.
5. Proof of Theorem 1.4
We prove Theorem 1.4 by following the strategy of [2]. In the above section, we have seen that for $i=1,2$, $P(\mathcal{A}_{n,i}^{c})$ decays exponentially fast. Here we show that $P\left(\mathcal{A}_{n,3}^{c}\right)$ also decays exponentially fast.
Recall the definition of $F_{m}$ in (2.1). For notational convenience, let $\widetilde{F}_{m}(t):=\exp\left(F_{m}(t)\right)$.
Lemma 5.1.
$\lim_{t\to\infty}\frac{\widetilde{F}_{m}(t)}{t^{2}}=1.$
Proof.
The statement is equivalent to
$\lim_{t\to\infty}F_{m}(t)-\log(1+t^{2})=0.$ (5.1)
We see that
$F_{m}(t)-\log(1+t^{2})=c_{m}\int_{\mathbb{R}}\frac{\log(1+(x-t)^{2})-\log(1+t^{2})}{(1+x^{2})^{m}}dx$
and
$\left|\log(1+(x-t)^{2})-\log(1+t^{2})\right|\leq\log(2(1+x^{2})).$
Now we can apply the Lebesgue convergence theorem. ∎
The following corresponds to [2, (3.15)].
Lemma 5.2.
Let $r>0$. Assume that
$0<\delta<\min\left\{\frac{1}{2}\left(m-\frac{1}{2}\right),\frac{F_{m}(r)-F_{m}(0)}{4}\right\}.$
Then, there exists a positive constant $c_{m,3}$ depending only on $m$ such that for every $t$ with $|t|>r$ and every $n\geq 1$,
$P(L_{n}(t)\leq F_{m}(0)+\delta)\leq\exp\left(-\frac{\delta}{c_{m,3}}(F_{m}(t)-F_{m}(0)-2\delta)n\right).$
Proof.
We assume that $t>r$. The proof is the same for the case that t<rt<-r. We see that
$P(L_{n}(t)\leq F_{m}(0)+\delta)=P\left(\sum_{i=1}^{n}(F_{m}(t)-\log(1+(X_{i}-t)^{2}))\geq n(F_{m}(t)-F_{m}(0)-\delta)\right).$
It holds that $F_{m}(t)-F_{m}(0)-\delta\geq F_{m}(r)-F_{m}(0)-\delta>0$ by Lemma 2.5. By the exponential Chebyshev inequality,
$P\left(\sum_{i=1}^{n}(F_{m}(t)-\log(1+(X_{i}-t)^{2}))\geq n(F_{m}(t)-F_{m}(0)-\delta)\right)$
$\leq\left(\exp\left(-\lambda(F_{m}(t)-F_{m}(0)-\delta)\right)E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]\right)^{n}$
for every $\lambda>0$.
Assume that $0<\lambda<m-1/2$. Then,
$E\left[\exp\left(\lambda\left|F_{m}(t)-\log(1+(X_{1}-t)^{2})\right|\right)\right]\leq\exp(\lambda F_{m}(t))E\left[(1+(X_{1}-t)^{2})^{\lambda}\right]<\infty.$
Therefore, we can apply the Taylor expansion and obtain that
$E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]=\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right].$
Since $E\left[\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right]=0$,
$\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]=1+\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right].$
By Lemma 5.1,
$c_{m,4}:=\sup_{x\leq t/2,t\geq 0}\log\frac{\widetilde{F}_{m}(t)}{1+(x-t)^{2}}<\infty.$
Hence,
$E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]\leq c_{m,4}^{k}P(X_{1}\leq t/2)+F_{m}(t)^{k}P(X_{1}>t/2).$
Since
$P(X_{1}>t/2)\leq c_{m}\int_{t/2}^{\infty}x^{-2m}dx\leq c_{m}4^{m}t^{1-2m},$
$E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]\leq c_{m,4}^{k}+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{k},$
where we let $c_{m,5}:=c_{m}4^{m}$. Hence,
$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]$
$\leq\frac{\lambda^{2}}{2}\left(c_{m,4}^{2}\exp(\lambda c_{m,4})+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda F_{m}(t))\right).$
Since $0<\lambda<m-\frac{1}{2}$,
$\lim_{t\to\infty}\frac{F_{m}(t)^{2}\exp(\lambda F_{m}(t))}{t^{2m-1}}=0,$
and hence,
$\sup_{t\geq 0}\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda F_{m}(t))<\infty.$
Let $\lambda_{m}:=\frac{1}{2}(m-\frac{1}{2})$. Then, for every $\lambda\in(0,\lambda_{m})$,
$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]\leq\lambda^{2}c_{m,6},$
where we let
$c_{m,6}:=\frac{1}{2}\left(c_{m,4}^{2}\exp(\lambda_{m}c_{m,4})+\sup_{t\geq 0}\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda_{m}F_{m}(t))\right)<\infty.$
We can assume that $c_{m,6}\geq 1$ because if $c_{m,6}<1$, then we can replace $c_{m,6}$ with $c_{m,6}+1$.
Therefore, for every $\lambda\in\left(0,\min\left\{\lambda_{m},\frac{F_{m}(r)-F_{m}(0)}{4}\right\}\right)$,
$E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]\leq\exp(\lambda^{2}c_{m,6}).$
If we let $\lambda:=\delta/c_{m,6}$, then, $0<\lambda<m-1/2$, and,
$\exp\left(-\lambda(F_{m}(t)-F_{m}(0)-\delta)\right)E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]$
$\leq\exp\left(-\frac{\delta}{c_{m,6}}(F_{m}(t)-F_{m}(0)-2\delta)\right).$
Thus, the assertion holds for $c_{m,3}=c_{m,6}$. ∎
Let $\lambda_{m}(r):=\frac{1}{2}\min\left\{\frac{1}{2}\left(m-\frac{1}{2}\right),\frac{F_{m}(r)-F_{m}(0)}{4}\right\}$.
The following corresponds to [2, (3.21)] [^4] There is a typo in [2, (3.21)]. The supremum in [2, (3.21)] should be the infimum..
Lemma 5.3.
Let $r>0$. Assume that $0<\delta<\lambda_{m}(r)$. Then, there exists $N(r,\delta)\in\mathbb{N}$ such that for every $n\geq N(r,\delta)$,
$P\left(\inf_{|t|\geq r}L_{n}(t)<F_{m}(0)+\delta\right)\leq 2\exp\left(-\frac{\delta^{2}}{8c_{m,3}}n\right),$
where $c_{m,3}$ is the constant appearing in Lemma 5.2.
We remark that $r>0$ can be taken arbitrarily small.
Proof.
We show that
$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)\leq\exp\left(-\frac{\delta^{2}}{8c_{m,3}}n\right).$ (5.2)
Since $L_{n}^{\prime}(t)=-2D_{n}(t)$ and $|D_{n}(t)|\leq 1/2$,
$\left\{\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right\}\subset\bigcup_{k\geq 1}\left\{L_{n}(k\delta+r)<F_{m}(0)+2\delta\right\}$
and hence, by Lemma 5.2,
$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta+r)<F_{m}(0)+2\delta\right)$
$\leq\sum_{k=1}^{\infty}\exp\left(-\frac{\delta}{c_{m,3}}(F_{m}\left(k\delta+r\right)-F_{m}(0)-4\delta)n\right)$
$\exp\left(-\frac{\delta}{c_{m,3}}(F_{m}(r)-F_{m}(0)-4\delta)n\right)\sum_{k=1}^{\infty}\exp\left(-\frac{\delta}{c_{m,3}}(F_{m}\left(k\delta+r\right)-F_{m}(r))n\right)$
$\leq\exp\left(-\frac{\delta^{2}}{8c_{m,3}}n\right)\sum_{k=1}^{\infty}\exp\left(-\frac{\delta}{c_{m,3}}(F_{m}\left(k\delta+r\right)-F_{m}(r))n\right).$
By (5.1), there exists a positive constant $T_{m,r}$ such that for every $t>T_{m,r}$, $F_{m}(t)\geq F_{m}(r)+\log t$. Hence, there exists $N_{T_{m,r}}\in\mathbb{N}$ such that for every $k>N_{T_{m,r}}$, $F_{m}(k\delta+r)\geq F_{m}(r)+\log(k\delta+r)$. Since
$\sum_{k=1}^{\infty}\exp\left(-\frac{\delta}{c_{m,3}}(F_{m}\left(k\delta+r\right)-F_{m}(r))n\right)$
$\leq N_{T_{m,r}}\exp\left(-\frac{\delta}{c_{m,3}}(F_{m}\left(\delta+r\right)-F_{m}(r))n\right)+\sum_{k=N_{T_{m,r}}+1}^{\infty}(k\delta+r)^{-n\delta/c_{m,3}}.$
Hence, for large $n$,
$\sum_{k=1}^{\infty}\exp\left(-\frac{\delta}{c_{m,3}}(F_{m}\left(k\delta+r\right)-F_{m}(r))n\right)\leq 1.$
Thus (5.2) holds.
The case that $t\leq-r$ can be dealt with in the same manner. ∎
The following corresponds to [2, (3.25)] [^5] There is also a typo in [2, (3.25)]. $n^{2}$” in the right hand side of the inequality in [2, (3.25)] should be $n\delta^{2}$”.. Recall that $\lambda_{m}=\frac{1}{2}(m-\frac{1}{2})$.
Lemma 5.4.
There exists a positive constant $c_{m,7}$ depending only on $m$ such that for every $\delta\in(0,c_{m,7}\lambda_{m})$ and every $n\geq 1$,
$P\left(L_{n}(0)\geq F_{m}(0)+\delta\right)\leq\exp\left(-\frac{n\delta^{2}}{2c_{m,7}}\right).$
Proof.
Assume that $0<\lambda\leq\lambda_{m}$. Then, by the exponential Chebyshev inequality,
$P(L_{n}(0)\geq F_{m}(0)+\delta)\leq\left(\exp(-\lambda\delta)E\left[\exp(\lambda(\log(1+X_{1}^{2})-F_{m}(0)))\right]\right)^{n}.$
Since $E[\log(1+X_{1}^{2})]=F_{m}(0)$,
$E\left[\exp(\lambda(\log(1+X_{1}^{2})-F_{m}(0)))\right]\leq\exp\left(\frac{\lambda^{2}}{2}c_{m,7}\right),$
where we let
$c_{m,7}:=E\left[(\log(1+X_{1}^{2})-F_{m}(0)))^{2}\exp\left(\lambda_{m}\left|\log(1+X_{1}^{2})-F_{m}(0))\right|\right)\right].$
Now let $\lambda:=\delta/c_{m,7}$. ∎
Let
$c_{m,8}:=\dfrac{1}{2}\min\left\{\lambda_{m}(r_{m,2}/3),c_{m,7}\lambda_{m}\right\}.$
Let $\mathcal{B}_{n,1}$ be the event that $\inf_{|t|\geq r_{m,2}/3}L_{n}(t)<F_{m}(0)+c_{m,8}$. Let $\mathcal{B}_{n,2}$ be the event that $L_{n}(0)\geq F_{m}(0)+c_{m,8}$. Then, $\hat{\theta}_{n}\in[-r_{m,2}/2,r_{m,2}/2]$ on the event $\mathcal{B}_{n,1}\cap\mathcal{B}_{n,2}$. Therefore,
$\mathcal{A}_{n,1}\cap\mathcal{A}_{n,2}\cap\mathcal{B}_{n,1}\cap\mathcal{B}_{n,2}\subset\mathcal{A}_{n}.$
By Lemma 3.1, Lemma 3.3, Lemma 5.3, and Lemma 5.4, there exist constants $c_{m,9},c_{m,10}$ depending only on $m$ such that for every $n\geq 1$,
$P(\mathcal{A}_{n}^{c})\leq P(\mathcal{A}_{n,1}^{c})+P(\mathcal{A}_{n,2}^{c})+P(\mathcal{B}_{n,1}^{c})+P(\mathcal{B}_{n,2}^{c})\leq c_{m,9}\exp(-c_{m,10}n).$
For $\epsilon\in(0,r_{m,2}/4)$,
$P(\{\hat{\theta}_{n}>\epsilon\}\cap\mathcal{A}_{n})=P(\{D_{n}(\epsilon)>0\}\cap\mathcal{A}_{n})$
and hence,
$\left|P\left(\hat{\theta}_{n}>\epsilon\right)-P(D_{n}(\epsilon)>0)\right|\leq 2P(\mathcal{A}_{n}^{c})\leq 2c_{m,9}\exp(-c_{m,10}n),\ n\geq 1.$ (5.3)
Let
$H_{m}(\epsilon):=\textup{Var}(D(X_{1},\epsilon))=E\left[D(X_{1},\epsilon)^{2}\right]-G_{m}(\epsilon)^{2}.$
Lemma $5.$
(1) $G_{m}(\epsilon)=-\frac{2m-1}{2(m+1)}\epsilon+O(\epsilon^{2}),\ \epsilon\to+0$.
(2) $H_{m}(\epsilon)=\frac{2m-1}{4m(m+1)}+O(\epsilon),\ \epsilon\to+0$.
Proof.
(1) By (3.8),
$\left|D(X_{1},\epsilon)-D(X_{1},0)-\epsilon\partial_{t}D(X_{1},0)\right|\leq C_{1}\epsilon^{2}.$ (5.4)
By (3.2) and (3),
$E[D(X_{1},0)]=0,E\left[\partial_{t}D(X_{1},0)\right]=-\frac{2m-1}{2(m+1)}.$
The estimate follows from these equalities and (5.4).
(2) By (5.4), there exists a positive constant $C_{2}$ such that for every $\epsilon\in(0,1)$,
$\left|D(X_{1},\epsilon)^{2}-D(X_{1},0)^{2}-2\epsilon D(X_{1},0)\partial_{t}D(X_{1},0)\right|\leq C_{2}\epsilon^{2}.$ (5.5)
Since $D(X_{1},0)$ and $\partial_{t}D(X_{1},0)$ are bounded, $D(X_{1},0)\partial_{t}D(X_{1},0)$ is also bounded, and in particular, is integrable. By (3.3),
$H_{m}(0)=E[D(X_{1},0)^{2}]=\frac{2m-1}{4m(m+1)}.$
The estimate follows from this equality and (5.5). ∎
We show (i). We consider the asymptotics of $P(D_{n}(\epsilon)>0)$.
We first give the upper estimate. We remark that $|D(X_{i},\epsilon)-G_{m}(\epsilon)|\leq\frac{1}{2}-G_{m}(\epsilon)$ and by Lemma 5.5,
$\lim_{\epsilon\to+0}G_{m}(\epsilon)\left(\frac{1}{2}-G_{m}(\epsilon)\right)=0,$
and,
$\lim_{\epsilon\to+0}H_{m}(\epsilon)=H_{m}(0)>0.$
Hence, there exists a constant $\epsilon_{m,1}>0$ depending only on $m$ such that for every $\epsilon\in(0,\epsilon_{m,1})$,
$\left|D(X_{i},\epsilon)-G_{m}(\epsilon)\right|\leq H_{m}(\epsilon).$
Lemma 5.6 (Petrov [11, Lemma 7.1] [^6] The statement is a little different from [2, Lemma 1]. In [2, Lemma 1], this assertion holds for large $n$, but this is valid for every $n\geq 1$.).
Let $Z_{i},i\geq 1$, be i.i.d. random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}:=\textup{Var}(Z_{1})>0$. Then, for every $n\geq 1$ and every $x\in[0,\sigma^{2}/M]$,
$P\left(\sum_{i=1}^{n}Z_{i}\geq nx\right)\leq\exp\left(-\frac{nx^{2}}{2\sigma^{2}}\left(1-\frac{Mx}{2\sigma^{2}}\right)\right).$
By this lemma, it holds that for every $\epsilon\in(0,\epsilon_{m,1})$ and every $n\geq 1$,
$P(D_{n}(\epsilon)>0)$ $=P\left(\sum_{i=1}^{n}D(X_{i},\epsilon)-G_{m}(\epsilon)>-nG_{m}(\epsilon)\right)$
$\leq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\right).$ (5.6)
By Lemma 5.5,
$\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\sim\frac{m(2m-1)}{m+1}\epsilon^{2},\ \epsilon\to+0,$ (5.7)
in particular,
$\lim_{\epsilon\to+0}\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)=0.$
By this, (5), and (5.3), it holds that there exists $\epsilon_{m,2}>0$ such that for every $\epsilon\in(0,\epsilon_{m,2})$, there exists $N_{\epsilon}$ such that for every $n\geq N_{\epsilon}$,
$P\left(\hat{\theta}_{n}>\epsilon\right)\leq 2\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\right).$
Hence, for every $\epsilon\in(0,\epsilon_{m,2})$,
$\limsup_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon\right)}{n}\leq-\frac{G_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right).$
By this, Lemma 5.5,
$\limsup_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\limsup_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon\right)}{n}\right)\leq-\frac{m(2m-1)}{2(m+1)}.$
The same argument is applicable to $P\left(\hat{\theta}_{n}<-\epsilon\right)$ and we obtain (1.1).
We next give the lower estimate. By Lemma 5.5,
$\lim_{\epsilon\to+0}G_{m}(\epsilon)=0\textup{ and }\lim_{\epsilon\to+0}H_{m}(\epsilon)=E\left[D(X_{1},0)^{2}\right]>0.$
Lemma 5.7 (Petrov [11, Lemma 7.2] [^7] The statement is a little different from [11, Lemma 7.2], however, we can show this assertion in the same manner as in the proof of [11, Lemma 7.2].).
Let $Z_{i},i\geq 1$, be i.i.d. random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}:=\textup{Var}(Z_{1})>0$. Then, for every $\eta>0$, there exists $r>0$ such that for every $x\in[0,r]$, there exists $N$ such that for every $n\geq N$,
$P\left(\sum_{i=1}^{n}Z_{i}\geq nx\right)\geq\exp\left(-\frac{nx^{2}}{2\sigma^{2}}\left(1+\eta\right)\right).$
By this lemma, for every $\eta>0$, there exists $\epsilon_{\eta}>0$ depending on $m$ and $\eta$ such that for every $\epsilon\in(0,\epsilon_{\eta})$, there exists $N_{\eta,\epsilon,1}\in\mathbb{N}$ such that for every $n\geq N_{\eta,\epsilon,1}$,
$P(D_{n}(\epsilon)>0)\geq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}(1+\eta)\right).$ (5.8)
In the same manner as in the upper bound, it holds that there exists $\epsilon_{\eta,2}>0$ depending on $\eta$ such that for every $\epsilon\in(0,\epsilon_{\eta,2})$, there exists $N_{\eta,\epsilon,2}\in\mathbb{N}$ such that for every $n\geq N_{\eta,\epsilon,2}$,
$P\left(\hat{\theta}_{n}>\epsilon\right)\geq\frac{1}{2}\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}(1+\eta)\right).$
Hence, for every $\epsilon\in(0,\epsilon_{\eta,2})$,
$\liminf_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon\right)}{n}\geq-\frac{G_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}(1+\eta).$
By this and Lemma 5.5, letting $\eta\to+0$,
$\liminf_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\liminf_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon\right)}{n}\right)\geq-\frac{m(2m-1)}{2(m+1)}.$
The same argument is applicable to $P\left(\hat{\theta}_{n}<-\epsilon\right)$ and we obtain (1.2).
Now we show (ii), but the proof is almost identical to the proof of (i).
By (5), it holds that for large $n$,
$P(D_{n}(\epsilon/\lambda_{n})>0)\leq\exp\left(-\frac{nG_{m}(\epsilon/\lambda_{n})^{2}}{2H_{m}(\epsilon/\lambda_{n})}\left(1+\frac{G_{m}(\epsilon/\lambda_{n})}{2H_{m}(\epsilon/\lambda_{n})}\right)\right).$
By Lemma 5.5,
$\lim_{n\to\infty}\lambda_{n}^{2}\frac{G_{m}(\epsilon/\lambda_{n})^{2}}{H_{m}(\epsilon/\lambda_{n})}\left(1+\frac{G_{m}(\epsilon/\lambda_{n})}{2H_{m}(\epsilon/\lambda_{n})}\right)=\frac{m(2m-1)}{m+1}.$
Therefore, we obtain that
$\limsup_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/\lambda_{n})>0\right)}{n/\lambda_{n}^{2}}\leq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$ (5.9)
By (5.8) and Lemma 5.5, we obtain that
$\liminf_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/\lambda_{n})>0\right)}{n/\lambda_{n}^{2}}\geq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$ (5.10)
(5.9) and (5.10) imply that
$\lim_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/\lambda_{n})>0\right)}{n/\lambda_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
By this and (5.3),
$\lim_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon/\lambda_{n}\right)}{n/\lambda_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
$P\left(\hat{\theta}_{n}<-\epsilon/\lambda_{n}\right)$ can be dealt with in the same manner.
Remark 5.8.
(i) Let $K(\cdot|\cdot)$ be the Kullback-Leibler divergence. Then, by computations,
$K\left(\textup{PVII}_{m}(\theta_{1},1)|\textup{PVII}_{m}(\theta_{2},1)\right)=m(F_{m}(\nu_{1}-\nu_{2})-F_{m}(0)).$
Let
$b(\epsilon,\theta):=\inf\left\{K\left(\textup{PVII}_{m}(\theta^{\prime},1)|\textup{PVII}_{m}(\theta,1)\right)\middle||\theta^{\prime}-\theta|>\epsilon\right\}.$
Since $F_{m}$ is symmetric and $t\mapsto F_{m}(|t|)$ is increasing, $b(\epsilon,\theta)=m(F_{m}(\epsilon)-F_{m}(0))$. Since $F_{m}^{\prime}=-2G_{m}$,
$\lim_{\epsilon\to+0}\frac{b(\epsilon,\theta)}{\epsilon^{2}}=\frac{m(2m-1)}{2(m+1)}=\frac{1}{I(\theta)}.$
(ii) For the case that $m=1$, the Bahadur efficiency for the joint estimation of the location and the scale is established in [1, Theorem 4] when both the location and the scale are unknown.
(iii) Gao [8] obtained moderate deviation results for the maximum likelihood estimator in a more general framework under certain regular conditions. Our model does not satisfy the conditions because the likelihood equation $D_{n}(t)=0$ has multiple roots.
6. Proof of Theorem 1.5
We deal with $P\left(\hat{\theta}_{n}>r\right)$. We see that for every $r>0$ and every $\delta>0$,
$P\left(\hat{\theta}_{n}>r\right)\leq P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)+P\left(L_{n}(0)\geq F_{m}(0)+\delta\right).$ (6.1)
We derive upper bounds of these probabilities by modifying the assertions in Section 5. The main difference is $\delta=\delta_{r}$ diverges when $r$ tends to infinity.
First, we give a lemma similar to Lemma 5.2. The proof differs in part. Recall that $\lambda_{m}=\frac{1}{2}(m-\frac{1}{2})$.
Lemma 6.1.
There exist two constants $r_{m,3}$ and $c_{m,11}$ such that for every $t\geq r_{m,3}$ and every $n\geq 1$,
$P\left(L_{n}(t)\leq F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\leq c_{m,11}t^{-n\lambda_{m}}.$
Proof.
As in Lemma 5.2, by the exponential Chebyshev inequality,
$P\left(L_{n}(t)\leq F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)$
$P\left(\sum_{i=1}^{n}(F_{m}(t)-\log(1+(X_{i}-t)^{2}))\geq n\left(F_{m}(t)-F_{m}(0)-\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)$
$\leq\left(\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}}\right]\right)^{n}.$
It holds that
$E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}}\right]$
$E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}},\ X_{1}\geq t/2\right]+E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}},\ X_{1}<t/2\right]$
$\leq P(X_{1}\geq t/2)+\left(1+\frac{t^{2}}{4}\right)^{-\lambda_{m}}=O(t^{1-2m}),\ \ t\to\infty.$
By (5.1),
$\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)=O\left(t^{8\lambda_{m}/3}\right),\ \ t\to\infty.$
Therefore,
$\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}}\right]=O(t^{-4\lambda_{m}/3}),\ t\to\infty.$
This completes the proof. ∎
Next, we give a lemma similar to Lemma 5.3. The proof is also similar.
Lemma 6.2.
There exist two positive constants $r_{m,4}$ and $c_{m,12}$ and $N_{m,1}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m,4}$ and every $n\geq N_{m,1}$,
$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\frac{F_{m}(r)-F_{m}(0)}{4}\right)\leq c_{m,12}r^{-n\lambda_{m}/2},$
Proof.
For notational convenience, let $\delta_{r}:=\dfrac{F_{m}(r)-F_{m}(0)}{4}$. Since $|L_{n}^{\prime}(t)|\leq 1$,
$\left\{\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{r}\right\}\subset\bigcup_{k\geq 1}\left\{L_{n}(k\delta_{r}+r)<F_{m}(0)+2\delta_{r}\right\}$
and hence, by Lemma 6.1, for every $n\geq 2/\lambda_{m}$ and every $r\geq r_{m,4}$,
$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{r}\right)\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{r}+r)<F_{m}(0)+2\delta_{r}\right)$
$\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{r}+r)<F_{m}(0)+\frac{F_{m}(k\delta_{r}+r)-F_{m}(0)}{2}\right)$
$\leq c_{m,11}\sum_{k=1}^{\infty}(k\delta_{r}+r)^{-n\lambda_{m}}\leq c_{m,11}\delta_{r}\int_{r}^{\infty}x^{-n\lambda_{m}}dx\leq c_{m,11}\delta_{r}r^{1-n\lambda_{m}}.$
By (5.1), $\delta_{r}=O(\log r),r\to\infty$ and we have the assertion. ∎
Finally, we give a lemma similar to Lemma 5.4.
Lemma 6.3.
There exist positive constants $r_{m,5}$ and $c_{m,13}$ such that for every $r>r_{m,5}$ and every $n\geq 1$,
$P\left(L_{n}(0)\geq F_{m}(0)+\delta_{r}\right)\leq r^{-nc_{m,13}}.$
Proof.
Let $c_{m,7}$ be the constant as in the proof of Lemma 5.4. Assume that $0<\lambda\leq\lambda_{m}$. Then, by the exponential Chebyshev inequality, for every $n\geq 1$,
$P(L_{n}(0)\geq F_{m}(0)+\delta_{r})\leq\left(\exp\left(-\lambda\delta_{r}+\frac{\lambda^{2}}{2}c_{m,7}\right)\right)^{n}.$
By (5.1), there exists a positive constant $r_{m,5}$ such that for every $r>r_{m,5}$, $2c_{m,7}\leq\log r\leq\delta_{r}$. Let $\lambda_{m}^{\prime}:=\min\{1,\lambda_{m}\}$. Thus, for every $r>r_{m,5}$ and every $n\geq 1$,
$P(L_{n}(0)\geq F_{m}(0)+\delta_{r})\leq\left(\exp\left(\left(-\lambda_{m}^{\prime}+\frac{(\lambda_{m}^{\prime})^{2}}{4}\right)\delta_{r}\right)\right)^{n}\leq r^{-nc_{m,13}},$
where we let $c_{m,13}:=\lambda_{m}^{\prime}-\frac{(\lambda_{m}^{\prime})^{2}}{4}>0$. ∎
By applying (6.1) to $\delta=\delta_{r}$, by Lemma 6.2 and Lemma 6.3, there exist positive constants $c_{m},r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
$P\left(\hat{\theta}_{n}>r\right)\leq r^{-c_{m}n}.$
$P\left(\hat{\theta}_{n}<-r\right)$ can be dealt with in the same manner, and we obtain Theorem 1.5.
7. Proof of Theorem 1.6
For $M>0$, let $\phi_{M}(x):=x^{2}\wedge M^{2}$. This is bounded and continuous on $\mathbb{R}$. By Theorem 1.2,
$\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}\phi_{M}(x)\varphi_{m}(x)dx,$ (7.1)
where $\varphi_{m}$ is the density function of the distribution $N\left(0,\frac{m+1}{m(2m-1)}\right)$.
Since $x^{2}\geq\phi_{M}(x)$,
$\liminf_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\geq\int_{\mathbb{R}}\phi_{M}(x)\varphi_{m}(x)dx.$
By the monotone convergence theorem,
$\liminf_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\geq\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx.$ (7.2)
We will show that
$\limsup_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\leq\frac{m+1}{m(2m-1)}.$ (7.3)
By (7.1) and the monotone convergence theorem,
$\lim_{M\to\infty}\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx.$
Hence, it suffices to show that
$\limsup_{M\to\infty}\left(\limsup_{n\to\infty}E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]\right)=0.$ (7.4)
By Fubini’s theorem for non-negative measurable functions and the change of variables $t=\sqrt{s}$,
$E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-M^{2},\ \left|\sqrt{n}\hat{\theta}_{n}\right|\geq M\right]$
$2\int_{M}^{\infty}tP\left(\left|\sqrt{n}\hat{\theta}_{n}\right|>t\right)dt=2n\int_{M/\sqrt{n}}^{\infty}sP\left(\left|\hat{\theta}_{n}\right|>s\right)ds.$
We consider $P\left(\hat{\theta}_{n}>s\right)$.
By (5.7), there exists $\epsilon_{m,3}\in(0,r_{m})$ such that for every $\epsilon\in(0,2\epsilon_{m,3})$,
$\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\geq\frac{m(2m-1)}{4(m+1)}\epsilon^{2}.$ (7.5)
Now we decompose the last integral into three parts:
$\int_{M/\sqrt{n}}^{\infty}=\int_{M/\sqrt{n}}^{\epsilon_{m,3}}+\int_{\epsilon_{m,3}}^{r_{m}+1}+\int_{r_{m}+1}^{\infty},$
where $r_{m}$ is the constant in Theorem 1.5.
By (7.5), (5), and (5.3), there exist two positive constants $c_{m,14},c_{m,15}$ and $N_{m,2}\in\mathbb{N}$ depending only on $m$ such that for every $n\geq N_{m,2}$ and $s\in(0,2\epsilon_{m,3})$,
$P\left(\hat{\theta}_{n}>s\right)\leq\exp\left(-\frac{m(2m-1)}{4(m+1)}s^{2}n\right)+c_{m,14}\exp(-c_{m,15}n).$
Therefore, for $n\geq N_{m,2}$,
$2n\int_{M/\sqrt{n}}^{\epsilon_{m,3}}sP\left(\hat{\theta}_{n}>s\right)ds$
$\leq\int_{M/\sqrt{n}}^{\epsilon_{m,3}}2ns\exp\left(-\frac{m(2m-1)}{4(m+1)}ns^{2}\right)ds+n\epsilon_{m,3}^{2}c_{m,14}\exp(-c_{m,15}n)$
$\leq\frac{4(m+1)}{m(2m-1)}\exp\left(-\frac{m(2m-1)}{4(m+1)}M^{2}\right)+n\epsilon_{m,3}^{2}c_{m,14}\exp(-c_{m,15}n).$
Hence,
$\limsup_{n\to\infty}2n\int_{M/\sqrt{n}}^{\epsilon_{m,3}}sP\left(\hat{\theta}_{n}>s\right)ds\leq\frac{4(m+1)}{m(2m-1)}\exp\left(-\frac{m(2m-1)}{4(m+1)}M^{2}\right).$ (7.6)
Since
$2n\int_{\epsilon_{m,3}}^{r_{m}+1}sP\left(\hat{\theta}_{n}>s\right)ds\leq 2(r_{m}+1)^{2}nP\left(\hat{\theta}_{n}>\epsilon_{m,3}\right),$
$\limsup_{n\to\infty}2n\int_{\epsilon_{m,3}}^{r_{m}+1}sP\left(\hat{\theta}_{n}>s\right)ds=0.$ (7.7)
By Theorem 1.5, for large $n$,
$n\int_{r_{m}+1}^{\infty}sP\left(\hat{\theta}_{n}>s\right)ds\leq n\int_{r_{m}+1}^{\infty}s^{1-nc_{m}}ds\leq\frac{n}{c_{m}n-2}(r_{m}+1)^{2-nc_{m}}.$
Hence,
$\limsup_{n\to\infty}2n\int_{r_{m}+1}^{\infty}sP\left(\hat{\theta}_{n}>s\right)ds=0.$ (7.8)
By (7.6), (7.7), and (7.8),
$\limsup_{n\to\infty}2n\int_{M/\sqrt{n}}^{\infty}sP\left(\hat{\theta}_{n}>s\right)ds\leq\frac{4(m+1)}{m(2m-1)}\exp\left(-\frac{m(2m-1)}{4(m+1)}M^{2}\right).$
The same estimate holds for $P\left(\hat{\theta}_{n}<-s\right)$. Since the right hand side converges to 0 as $M\to\infty$, (7.4) holds.
Thus we obtain (7.2) and (7.3) and the proof is completed.
Remark 7.1.
The variance of the maximum likelihood estimator of the parameter $m$ was dealt with by Taylor’s unpublished manuscript [14].
8. Numerical computations
We perform simulation studies using the software R to illustrate the properties of the maximum likelihood estimator. We used R version 4.5.1. We deal with $nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$ appearing in Theorem 1.6. We can assume that $\theta=0$ without loss of generality. We consider parameters $m=0.1\times k$ for $6\leq k\leq 15$ and sample sizes $n=10,50,100,500,1000$. In each choice of the pair $(m,n)$, we compute the average of $n\left(\hat{\theta}_{n}\right)^{2}$ over $10^{7}$ samples of size $n$ generated by the rpearsonVII() function in the package ‘PearsonDS’, and in the optimization, we use the nlminb() function with the starting point being the median. Table 1 gives the result.
In the case of $m=1$, [6, Table 2] gives numerical computations for $n=5,6,\dots,14,15,20,25,\dots,35,40$. This is consistent with the numerical computation in [1, Table 2] for the joint estimation of the location and scale.
The convergence becomes faster as the parameter $m$ increases. Since $\hat{\theta}_{n}$ appears not to be square-integrable if $m=0.6$ and $n=10$, and the numerical computation is not stable although $\hat{\theta}_{n}$ appears to be square-integrable if $m=0.7$ and $n=10$, we mark them as not available (NA). By this table, we conjecture that for each $m$, $\left(nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]\right)_{n}$ is decreasing in $n$.
[…]
0.6 NA 25.756 16.108 13.756 13.545 13.333
0.7 NA 7.236 6.566 6.160 6.116 6.071
0.8 9.046 4.156 3.935 3.783 3.766 3.750
0.9 4.524 2.832 2.729 2.655 2.648 2.639
1 2.908 2.108 2.052 2.011 2.006 2.000
1.1 2.103 1.659 1.624 1.599 1.595 1.591
1.2 1.630 1.356 1.332 1.314 1.311 1.310
1.3 1.320 1.138 1.123 1.108 1.107 1.106
1.4 1.105 0.977 0.964 0.954 0.953 0.952
1.5 0.945 0.851 0.842 0.835 0.834 0.833
Table 1. Simulated values of $n\,\mathbb{E}\!\left[(\hat{\theta}_{n}-\theta)^{2}\right]$ ($\theta=0$). Rows correspond to $m$; columns to the sample size $n$. The column labeled $\infty$ reports the theoretical limit (m+1)/(m(2m1))(m+1)/(m(2m-1)) given by Theorem 1.6. We round the results to three decimal places.

### v2

@@
 Key words and phrases:
-Pearson Type VII distribution, Cauchy distribution, maximum likelihood estimator, strong consistency, asymptotic efficiency, Bahadur efficiency
+Pearson Type VII distribution, Cauchy distribution, maximum likelihood estimator, strong consistency, asymptotic normality, asymptotic efficiency, Bahadur efficiency
 2020 Mathematics Subject Classification:
@@
 The family of Pearson Type VII distributions provides flexible heavy-tailed models.
-The estimation of its parameters dates back at least to Fisher [7], nearly a century ago, and many researchers have studied it since then; see Johnson, Kotz, and Balakrishnan [9, Section 28] for a thorough survey of results prior to 1994.
+The estimation of its parameters dates back at least to Fisher [10], over a century ago, and many researchers have studied it since then; see Johnson, Kotz, and Balakrishnan [14, Section 28] for a thorough survey of results prior to 1994.
 This class is also known as the location–scale family of Student’s $t$ distributions or of $q$-Gaussian distributions.
-For estimating the location, the median is a robust alternative to the arithmetic mean; however, it is not asymptotically efficient.
+For estimating the location, the median is a robust alternative to the arithmetic mean; however it is not asymptotically efficient.
 In general, the maximum likelihood estimator is widely regarded as optimal in large samples under standard regularity.
 Lange, Little, and Taylor [15] proposed a strategy based on maximum likelihood for a general model with errors following the $t$-distribution and applied it to many problems.
-Under suitable regularity conditions, properties such as strong consistency, asymptotic efficiency, and Bahadur efficiency have been established by many researchers.
+Under suitable regularity conditions, properties such as strong consistency, asymptotic normality, and Bahadur efficiency have been established by many researchers.
 For location–scale families, it is natural to consider the estimation of the location with known scale.
@@
 For this reason, alternative estimators of the Cauchy location parameter have been considered.
-For example, Cohen Freue [6] considered the Pitman estimator for small samples, and Zhang [17] considered an empirical Bayes estimator.
+For example, Freue [11] considered the Pitman estimator for small samples, and Zhang [24] considered an empirical Bayes estimator.
 Nevertheless, this does not represent a failure of the maximum likelihood estimator itself.
@@
 We show that the asymptotic properties of the maximum likelihood estimator mirror those for the arithmetic mean of independent and identically distributed (i.i.d.) random variables with finite variance.
-Asymptotically, the maximum likelihood estimator for the Pearson Type VII distribution behaves well.
+Asymptotically, the maximum likelihood estimator for the Pearson Type VII distribution performs well.
 Now we state the framework and the main result.
@@
 Let $\textup{PVII}_{m}(\mu,\sigma)$ be the Pearson Type VII distribution with location $\mu\in\mathbb{R}$ and scale $\sigma>0$.
-Then, the probability density function of $\textup{PVII}_{m}(\mu,\sigma)$ is given by
+Then the probability density function of $\textup{PVII}_{m}(\mu,\sigma)$ is given by
 $f(x)=c_{m}\frac{1}{\sigma}\left(1+\left(\frac{x-\mu}{\sigma}\right)^{2}\right)^{-m},$
-where $c_{m}$ is the normalizing constant, specifically, $c_{m}:=\left(\int_{\mathbb{R}}(1+x^{2})^{-m}dx\right)^{-1}$. The case that $m=1$ is the Cauchy distribution.
+where $c_{m}$ is the normalizing constant, specifically, $\displaystyle c_{m}\coloneqq\left(\int_{\mathbb{R}}(1+x^{2})^{-m}dx\right)^{-1}$. The case $m=1$ corresponds to the Cauchy distribution.
 We consider the maximum likelihood estimator of the location parameter of the Pearson Type VII distribution with known scale.
@@
 Let $(X_{n})_{n\geq 1}$ be i.i.d.
-random variables on a probability space $(\Omega,\mathcal{F},P)$ following $\textup{PVII}_{m}(\theta,1)$.
+random variables on a complete probability space $(\Omega,\mathcal{F},P)$ following $\textup{PVII}_{m}(\theta,1)$.
 Let $\hat{\theta}_{n}$ be the maximum likelihood estimator of the location parameter from a sample $(X_{1},\dots,X_{n})$ of size $n$.
-Let $\hat{\theta}_{n}(x_{1},\dots,x_{n})$ be a measurable function on $\mathbb{R}^{n}$ which maximizes the function $\theta\mapsto\prod_{i=1}^{n}f(x_{i}-\theta)$.
+Let $\hat{\theta}_{n}(x_{1},\dots,x_{n})$ be a measurable function on $\mathbb{R}^{n}$ which maximizes the function $\displaystyle\theta\mapsto\prod_{i=1}^{n}f(x_{i}-\theta)$.
 Such a function exists by virtue of the measurable selection theorem.
-Then, let $\hat{\theta}_{n}:=\hat{\theta}_{n}(X_{1},\dots,X_{n})$.
+Let $\hat{\theta}_{n}\coloneqq\hat{\theta}_{n}(X_{1},\dots,X_{n})$.
 Our first main result is strong consistency.
 Theorem 1.1 (Strong consistency).
-$\lim_{n\to\infty}\hat{\theta}_{n}=\theta,\textup{ $P$-a.s.}$
+$\lim_{n\to\infty}\hat{\theta}_{n}=\theta,\textup{ $\displaystyle P$-a.s.}$
 We show this using the concept of the Fréchet mean.
-Once we see the strong consistency, it is natural to consider the asymptotic normality. We denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$ by $N(\mu,\sigma^{2})$.
+Once the strong consistency is given, it is natural to consider the asymptotic normality. We denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$ by $N(\mu,\sigma^{2})$.
 Theorem 1.2 (Asymptotic normality).
-$\left(\sqrt{n}(\hat{\theta}_{n}-\theta)\right)_{n}$ converges to $N\left(0,\dfrac{m+1}{m(2m-1)}\right)$ in distribution as $n\to\infty$.
+$\left(\sqrt{n}(\hat{\theta}_{n}-\theta)\right)_{n}$ converges to
+$N\left(0,\dfrac{m+1}{m(2m-1)}\right)$ in distribution as $n\to\infty$.
 By Remark 3.5 below, $I(\theta)=\dfrac{m(2m-1)}{m+1}$, where $I(\theta)$ is the Fisher information for a single observation.
-We can also see the following behavior, although it is not often considered in mathematical statistics.
+We proceed to the law of the iterated logarithm. It has connections with statistics, in particular with sequential testing. See [18, 6, 13].
 Theorem 1.3 (Law of the iterated logarithm).
-$\limsup_{n\to\infty}\sqrt{\frac{n}{\log\log n}}(\hat{\theta}_{n}-\theta)=\sqrt{\frac{2(m+1)}{m(2m-1)}},\textup{ $P$-a.s.}$
+$\limsup_{n\to\infty}\sqrt{\frac{n}{\log\log n}}(\hat{\theta}_{n}-\theta)=\sqrt{\frac{2(m+1)}{m(2m-1)}},\textup{ $\displaystyle P$-a.s.}$
 For the proof, we use the technique of the deviation mean of i.i.d. random variables investigated by Barczy and Páles [4] with some modifications.
@@
 Theorem 1.4 (Bahadur efficiency and moderate deviation).
-$\limsup_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\limsup_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\leq-\frac{m(2m-1)}{2(m+1)}.$ (1.1)
-$\liminf_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\liminf_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\geq-\frac{m(2m-1)}{2(m+1)}.$ (1.2)
-(ii) For every sequence $(\lambda_{n})_{n}$ of positive numbers satisfying $\lim_{n\to\infty}\lambda_{n}=\infty$ and $\lim_{n\to\infty}\lambda_{n}/n^{1/2}=0$ and every $\epsilon>0$,
-$\lim_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon/\lambda_{n}\right)}{n/\lambda_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
-This assertion implies Theorem 1.1 and its proof does not depend on Theorem 1.1.
-However, we can show Theorem 1.1 much more easily than the proof of this assertion.
+(1.1) $\limsup_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\limsup_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\leq-\frac{m(2m-1)}{2(m+1)}.$
+(1.2) $\liminf_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\liminf_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon\right)}{n}\right)\geq-\frac{m(2m-1)}{2(m+1)}.$
+(ii) For every sequence $(a_{n})_{n}$ of positive numbers satisfying $\displaystyle\lim_{n\to\infty}a_{n}=\infty$ and $\displaystyle\lim_{n\to\infty}a_{n}/n^{1/2}=0$ and every $\epsilon>0$,
+$\lim_{n\to\infty}\frac{\log P\left(\left|\hat{\theta}_{n}-\theta\right|>\epsilon/a_{n}\right)}{n/a_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
+This assertion implies Theorem 1.1 and its proof does not depend on Theorem $1.$ However we can show Theorem 1.1 much more easily than the proof of this assertion.
 For the proof, we follow the strategy of [3].
-It is worth investigating the probability that the estimator deviates significantly from the true value. Hereafter, $\mathbb{N}=\{1,2,\dots\}$.
+It is worth investigating the probability that the estimator deviates significantly from the true value. In this paper, we let $\mathbb{N}\coloneqq\{1,2,\dots\}$.
 Theorem 1.5 (Integrability).
-There exist positive constants $c_{m},r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
-$P\left(\left|\hat{\theta}_{n}-\theta\right|>r\right)\leq r^{-c_{m}n}.$
-In particular, $\hat{\theta}_{n}\in L^{c_{m}n-1}(\Omega,\mathcal{F},P)$ for $n\geq N_{m}$.
+There exist positive constants $r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
+$P\left(\left|\hat{\theta}_{n}-\theta\right|>r\right)\leq r^{-c_{m}^{\prime}n},$
+where $c_{m}^{\prime}\coloneqq\lambda_{m}^{\prime}-\dfrac{(\lambda_{m}^{\prime})^{2}}{4}$ and $\lambda_{m}^{\prime}\coloneqq\min\left\{1,\dfrac{2m-1}{4}\right\}$.
+In particular, $\hat{\theta}_{n}\in L^{c_{m}^{\prime}n-1}(\Omega,\mathcal{F},P)$ for $n\geq N_{m}$.
 We show this by modifying several estimates in the proof of Theorem 1.4.
@@
 $\lim_{n\to\infty}nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]=\frac{m+1}{m(2m-1)}.$
-This is consistent with [9, (28.61c)]. We give a mathematically rigorous proof of it. The proof is technically involved and we use Theorems 1.2 and 1.5.
-In the following sections, we present proofs of these assertions. In the final section, we give numerical computations of $nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$.
+This is consistent with [14, (28.61c)].
+We give a mathematically rigorous proof of it.
+The proof is technically involved and we use Theorem 1.2 and Theorem 1.5, and an estimate obtained in the proof of Theorem 1.4.
+In the following sections, we present proofs of these assertions. In the final section, we give simulation studies.
 In the proofs of these results, we can assume that $\theta=0$ without loss of generality.
@@
 Many constants will appear.
-When a constant depends only on $m$, we indicate this by attaching $m$ to its index; otherwise we omit it even if it depends on $m$.
+When a constant depends only on $m$, we indicate this by including $m$ as a subscript; otherwise we omit it even if it depends on $m$.
 2. Proof of Theorem 1.1
-We prove Theorem 1.1 by following the strategy of [4, Section 3.2].
-One of our goals is to establish [4, Theorem 3.3] in the case where the loss function is replaced [^1] In [4, Theorem 3.3], the loss function is given by the map $u\mapsto u^{\alpha}$ for $\alpha\in(0,1)$.
+We first give an outline of the proof.
+We prove Theorem 1.1 by following the strategy of Bhattacharya and Bhattacharya [5, Section 3.2].
+One of our goals is to establish [5, Theorem 3.3] in the case where the loss function is replaced [^1] In [5, Theorem 3.3], the loss function is given by the map $u\mapsto u^{\alpha}$ for $\alpha\in(0,1)$.
 with the map $u\mapsto\log(1+u^{2})$.
+The key step is Proposition 2.7, which shows that the Fréchet mean set $C_{Q_{n}(\omega)}$ (equivalently, the argmin set of $L_{n}$) is eventually contained in an arbitrarily small neighborhood of $0$.
+Proposition 2.7 immediately yields Theorem $1.$ The proof of Proposition 2.7 consists of four ingredients: uniform boundedness of minimizers (Lemma 2.2), a uniform law of large numbers for $L_{n}$ on bounded intervals (Lemma 2.4), uniqueness of the population minimizer $C_{\nu_{m}}=\{0\}$ (Lemma 2.5), and a stability lemma for minimizers of continuous functions with compact level sets (Lemma 2.6).
 Let
-$L_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}\log(1+(X_{i}-t)^{2}),\ t\in\mathbb{R}.$
-Let $\nu_{m}$ be the Borel probability measure of the Pearson Type VII distribution $\textup{PVII}_{m}(0,1)$, that is,
-$\nu_{m}(dx)=c_{m}\left(1+x^{2}\right)^{-m}dx.$
+$L_{n}(t)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\log(1+(X_{i}-t)^{2}),\ t\in\mathbb{R}.$
+Let $\nu_{m}$ be the probability measure on the Lebesgue measurable space $(\mathbb{R},\mathcal{L}(\mathbb{R}))$ of the Pearson Type VII distribution $\textup{PVII}_{m}(0,1)$, that is,
+$\nu_{m}(dx)\coloneqq c_{m}\left(1+x^{2}\right)^{-m}dx.$
 Lemma 2.1.
@@
 By the strong law of large numbers,
-$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}{\bf 1}_{[-1,1]}(X_{i}(\omega))=\nu_{m}([-1,1])>0,\ \textup{ $P$-a.s.$\omega$}.$
-We have the assertion for $c_{m,1}:=\nu_{m}([-1,1])$. ∎
+$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}{\bf 1}_{[-1,1]}(X_{i}(\omega))=\nu_{m}([-1,1])>0,\ \textup{ $\displaystyle P$-a.s. $\displaystyle\omega$}.$
+We have the assertion for $c_{m,1}\coloneqq\nu_{m}([-1,1])$. ∎
 Denote the empirical distribution of $(X_{i}(\omega))_{i=1}^{n}$ by $Q_{n}(\omega)$, specifically,
-$Q_{n}(\omega)=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}(w)}.$
+$Q_{n}(\omega)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}(w)}.$
+For a probability measure $\nu$ on the Lebesgue measurable space $(\mathbb{R},\mathcal{L}(\mathbb{R}))$, let the expected loss function be
+$F_{\nu}(t)\coloneqq\int_{\mathbb{R}}\log(1+(x-t)^{2})d\nu(x),\ t\in\mathbb{R},$
+and the mean set be
+$C_{\nu}\coloneqq\left\{t\in\mathbb{R}\middle|\min_{s\in\mathbb{R}}F_{\nu}(s)=F_{\nu}(t)\right\}.$
+Since $\log(1+(x-t)^{2})=O\left(|x|^{\alpha}\right),x\to\infty$, for every $\alpha>0$, $F_{\nu_{m}}(t)<\infty$ for every $t\in\mathbb{R}$.
+For $\nu=Q_{n}(\omega)$, $C_{Q_{n}(\omega)}$ is called the Fréchet mean set.
+Recall that maximizing the likelihood is equivalent to minimizing the empirical negative log-likelihood $L_{n}$.
+For the empirical measure $Q_{n}$, the corresponding Fréchet function $F_{Q_{n}(\omega)}(t)$ equals $L_{n}(t)(\omega)$.
+Therefore, $\hat{\theta}_{n}(\omega)\in C_{Q_{n}(\omega)}$.
+Another goal of this section is to show that the mean set $C_{\nu_{m}}$ is a singleton, which will be established in Lemma 2.5 below. Let
+(2.1) $F_{m}(t)\coloneqq F_{\nu_{m}}(t)=\int_{\mathbb{R}}\log\left(1+(x-t)^{2}\right)\nu_{m}(dx).$
+This is the expected loss function.
 Lemma 2.2 (boundedness of minimizers).
@@
 $L_{n}(0)(\omega)\geq L_{n}(t_{n})(\omega)\geq\frac{c_{m,1}}{4}(\log(1+t_{n}^{2})-2\log 2)$
+Recall (2.1).
+We remark that $F_{\nu_{m}}(0)<\infty$.
 By the strong law of large numbers, there exists an event $\Omega_{2}\subset\Omega_{1}$ such that $P(\Omega_{2})=1$ and for every $\omega\in\Omega_{2}$,
@@
 Hence, there exists a constant $r_{m,1}>1$ such that for every $\omega\in\Omega_{2}$ and $n>N_{2}(\omega)$, $|t_{n}|<r_{m,1}$. ∎
-For a Borel probability measure $\nu$ on $\mathbb{R}$, let the expected loss function be
-$F_{\nu}(t):=\int_{\mathbb{R}}\log(1+(x-t)^{2})d\nu(x),\ t\in\mathbb{R},$
-and the mean set be
-$C_{\nu}:=\left\{t\in\mathbb{R}\middle|\min_{s\in\mathbb{R}}F_{\nu}(s)=F_{\nu}(t)\right\}.$
-For $\nu=Q_{n}(\omega)$, $C_{Q_{n}(\omega)}$ is called the Fréchet mean set. It holds that $\hat{\theta}_{n}(\omega)\in C_{Q_{n}(\omega)}$.
-Another goal of this section is to show that the mean set $C_{\nu_{m}}$ is a singleton, which will be established in Lemma 2.5 below.
-For notational convenience,
-$F_{m}(t):=F_{\nu_{m}}(t)=\int_{\mathbb{R}}\log(1+(x-t)^{2})\nu_{m}(dx).$ (2.1)
-This is the expected loss function.
 Lemma 2.3 (a.s. pointwise convergence).
-$P$-a.s., it holds that every $t\in\mathbb{R}$,
-$\lim_{n\to\infty}L_{n}(t)=F_{m}(t).$ (2.2)
+$P$-a.s., it holds that for every $t\in\mathbb{R}$,
+(2.2) $\lim_{n\to\infty}L_{n}(t)=F_{m}(t).$
 Proof.
-By the strong law of large numbers, for every fixed $t\in\mathbb{R}$, (2.2) holds a.s.
-Now, use the rational approximation and use the Lipschitz continuity [^2] This estimate works well if $x$ and $y$ are close.
-If $x$ or $y$ is large, the bound can be very loose.
-of $\log(1+x^{2})$, specifically,
-$\left|\log(1+x^{2})-\log(1+y^{2})\right|\leq\left||x|-|y|\right|\leq|x-y|.$ (2.3)
+By the strong law of large numbers, for every fixed $t\in\mathbb{R}$, equation (2.2) holds a.s.
+More specifically, for every $t\in\mathbb{R}$, there exists an event $\Omega_{t}$ such that $P(\Omega_{t})=1$ and for every $\omega\in\Omega_{t}$, $\lim_{n\to\infty}L_{n}(t)(\omega)=F_{m}(t)$.
+We use the Lipschitz continuity [^2] This estimate works well if $x$ and $y$ are close. If $x$ or $y$ is large, the bound can be very loose. of $\log(1+x^{2})$, specifically,
+(2.3) $\left|\log(1+x^{2})-\log(1+y^{2})\right|\leq\left||x|-|y|\right|\leq|x-y|.$
+Hence
+(2.4) $\left|L_{n}(t)(\omega)-L_{n}(s)(\omega)\right|\leq|t-s|,\ \ t,s\in\mathbb{R},\,n\geq 1,$
+and
+(2.5) $\left|F_{m}(t)-F_{m}(s)\right|\leq|t-s|,\ t,s\in\mathbb{R}.$
+We use the rational approximation.
+Let $\displaystyle\Omega_{\mathbb{Q}}\coloneqq\bigcap_{t\in\mathbb{Q}}\Omega_{t}$.
+Then $P(\Omega_{\mathbb{Q}})=1$.
+Take $t\in\mathbb{R}$ arbitrarily.
+By (2.4) and (2.5), it holds that for every $\omega\in\Omega_{\mathbb{Q}}$ and $s\in\mathbb{Q}$,
+$\limsup_{n\to\infty}|L_{n}(t)(\omega)-F_{m}(t)|\leq 2|t-s|+\limsup_{n\to\infty}|L_{n}(s)(\omega)-F_{m}(s)|=2|t-s|.$
+Since $s\in\mathbb{Q}$ can be taken arbitrarily close to $t$, we see that $\lim_{n\to\infty}|L_{n}(t)(\omega)-F_{m}(t)|=0$. ∎
 The following is the uniform law of large numbers.
@@
 Lemma 2.5.
-$F_{m}$ is strictly decreasing on $(-\infty,0)$ and strictly increasing on $(0,\infty)$. In particular, $C_{\nu_{m}}=\{0\}$.
+The function $F_{m}$ is strictly decreasing on $(-\infty,0)$ and strictly increasing on $(0,\infty)$. In particular, $C_{\nu_{m}}=\{0\}$.
 Proof.
 By the Lebesgue convergence theorem, we see that
-$F_{m}^{\prime}(t)=-2c_{m}\int_{\mathbb{R}}\frac{x-t}{(1+(x-t)^{2})(1+x^{2})^{m}}dx.$ (2.4)
+(2.6) $F_{m}^{\prime}(t)=-2c_{m}\int_{\mathbb{R}}\frac{x-t}{(1+(x-t)^{2})(1+x^{2})^{m}}dx.$
 By change of variables,
@@
 For a non-empty subset $A$ of $\mathbb{R}$, let
-$d(x,A):=\inf\left\{|x-y|:y\in A\right\}.$
+$d(x,A)\coloneqq\inf\left\{|x-y|:y\in A\right\}.$
 Lemma 2.6.
-Let $\varphi$ be a continuous function on $\mathbb{R}$ such that $\lim_{|z|\to\infty}\varphi(z)=\infty$.
-Let $C_{\varphi}:=\left\{x\in\mathbb{R}:\min_{t\in\mathbb{R}}\varphi(t)=\varphi(x)\right\}$.
-Then, for every $\epsilon>0$, there exists $\delta>0$ such that for every $x\in\mathbb{R}$ with $\varphi(x)\leq\min_{t\in\mathbb{R}}\varphi(t)+\delta$, $d(x,C_{\varphi})<\epsilon$.
+Let $\varphi$ be a continuous function on $\mathbb{R}$ such that $\displaystyle\lim_{|z|\to\infty}\varphi(z)=\infty$.
+Let $\displaystyle C_{\varphi}\coloneqq\left\{x\in\mathbb{R}:\min_{t\in\mathbb{R}}\varphi(t)=\varphi(x)\right\}$.
+Then, for every $\epsilon>0$, there exists $\delta>0$ such that for every $x\in\mathbb{R}$ with $\displaystyle\varphi(x)\leq\min_{t\in\mathbb{R}}\varphi(t)+\delta$, $d(x,C_{\varphi})<\epsilon$.
 Proof.
@@
 Assume that there exists $\epsilon_{0}>0$ such that for every $n\in\mathbb{N}$, there exists $x_{n}\in\mathbb{R}$ such that $\varphi(x_{n})\leq\min_{t\in\mathbb{R}}\varphi(t)+1/n$ and $d(x_{n},C_{\varphi})\geq\epsilon_{0}$.
-Since $\sup_{n\in\mathbb{N}}\varphi(x_{n})<+\infty$, by the assumption of $\varphi$, $(x_{n})_{n}$ is a bounded sequence.
-Then, there exist a subsequence $(x_{n_{k}})_{k}$ and a point $z\in\mathbb{R}$ such that $x_{n_{k}}\to z,k\to\infty$.
-By the continuity of $\varphi$, $\varphi(z)=\lim_{k\to\infty}\varphi(x_{n_{k}})=\min_{t\in\mathbb{R}}\varphi(t)$.
+Since $\displaystyle\sup_{n\in\mathbb{N}}\varphi(x_{n})<+\infty$, by the assumption of $\varphi$, $(x_{n})_{n}$ is a bounded sequence.
+Then there exist a subsequence $(x_{n_{k}})_{k}$ and a point $z\in\mathbb{R}$ such that $x_{n_{k}}\to z,k\to\infty$.
+By the continuity of $\varphi$, $\displaystyle\varphi(z)=\lim_{k\to\infty}\varphi(x_{n_{k}})=\min_{t\in\mathbb{R}}\varphi(t)$.
 Hence, $z\in C_{\varphi}$.
@@
 $\lim_{n\to\infty}\left|L_{n}(t_{n})(\omega)-F_{m}(t_{n})\right|=0.$
-Let $\epsilon>0$. Then, there exists $N_{4}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n>N_{4}(\omega,\epsilon)$ and $t_{n}\in C_{Q_{n}(\omega)}$,
+Let $\epsilon>0$. Then there exists $N_{4}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n>N_{4}(\omega,\epsilon)$ and $t_{n}\in C_{Q_{n}(\omega)}$,
 $F_{m}(t_{n})\leq L_{n}(t_{n})(\omega)+\frac{\epsilon}{4}$
@@
 $L_{n}(t_{n})\leq F_{m}(0)+\epsilon.$
-Now apply Lemma 2.6 to $\varphi=F_{m}$ and Lemma 2.5. ∎
-By Proposition 2.7, we obtain Theorem 1.1.
+Now apply Lemma 2.6 to $\displaystyle\varphi=F_{m}$ and Lemma 2.5. ∎
+Recall that $\hat{\theta}_{n}$ is a measurable selection from $C_{Q_{n}(\omega)}$. By Proposition 2.7, we obtain Theorem $1.$
 Remark 2.8.
-Recently, Schötz [13] gave precise analysis for the Fréchet mean. His approach uses a general ergodic theorem and differs from the above approach.
+Recently, Schötz [19] gave a precise analysis for the Fréchet mean. His approach uses a general ergodic theorem and differs from the above approach.
 3. Proof of Theorem 1.2
-We follow the strategy of [3, Section 4]. Let
-$D(x,t):=\frac{x-t}{1+(x-t)^{2}},\ x,t\in\mathbb{R}.$
+We first give an outline of the proof. We follow the strategy of Barczy and Páles [4, Section 4]. Let
+$D(x,t)\coloneqq\frac{x-t}{1+(x-t)^{2}},\ x,t\in\mathbb{R},$
+and
+$D_{n}(t)\coloneqq\frac{1}{n}\sum_{i=1}^{n}D(X_{i},t),\ t\in\mathbb{R}.$
+Then $-2D_{n}(t)\equiv L_{n}^{\prime}(t)$ and hence, the likelihood equation is $D_{n}(t)=0$.
+Since the map $t\mapsto D(x,t)$ is not monotone on $\mathbb{R}$ for each fixed $x$, we cannot apply the result of [4, Section 4] directly.
+Therefore we “localize” the argument.
+Specifically, we construct a sequence of events $\mathcal{A}_{n},n\geq 1$, defined in (3.2) below, such that $\displaystyle\lim_{n\to\infty}P(\mathcal{A}_{n})=1$, and, on each $\mathcal{A}_{n}$ the map $t\mapsto D_{n}(t)$ is strictly decreasing on a fixed interval $I$ and has a unique zero in $I$.
+This yields that on each $\mathcal{A}_{n}$, $\hat{\theta}_{n}>t$ if and only if $D_{n}(t)>0$ for $t\in I$.
+Once this localization is established, the remainder of the proof is standard.
+Consider a Taylor expansion of $D_{n}(y/\sqrt{n})$ at $0$ and apply the central limit theorem, the law of large numbers and Slutsky’s lemma.
+The arguments in Section 2 are not sufficient for this localization, since the possibility that $|C_{Q_{n}}\cap[-T,T]|\geq 2$ has not yet been excluded.
+We overcome this issue by Proposition 3.4.
+For the proof, we compare $D_{n}(t)$ and its derivative $D_{n}^{\prime}(t)$ with their population counterparts $G_{m}(t)$ and $G_{m}^{\prime}(t)$ defined below.
+Lemma 3.1 shows that $D_{n}(t)$ is uniformly close to $G_{m}(t)$ for large $n$.
+Lemma 3.2 shows that $G_{m}^{\prime}(t)<0$.
+Lemma 3.3 transfers this to $D_{n}^{\prime}(t)$ for large $n$.
+Theorem 1.2 can also be shown by using van der Vaart [22, Theorem 5.23].
+See Remark 3.5 (iii) below for details.
+However, throughout this paper we repeatedly use the notation such as $\{\mathcal{A}_{n}\}_{n\geq 1}$ and refer to related assertions in this section, we therefore present the full details here.
 Let
-$D_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}D(X_{i},t).$
-Then, $2D_{n}(t)\equiv S_{n}^{\prime}(t)$ and hence, the likelihood equation is $D_{n}(t)=0$.
-We first show that $\hat{\theta}_{n}>t$ holds if and only if $D_{n}(t)>0$ on an event with high probability.
-Since $D(x,t)$ is not monotone with respect to $t$ on $\mathbb{R}$, we cannot apply the result of [3] directly and need to “localize” it.
-The arguments in Section 2 are not sufficient for this goal, since the possibility that $|C_{Q_{n}}\cap[-T,T]|\geq 2$ has not yet been excluded.
-Let
-$G_{m}(t):=E\left[D(X_{1},t)\right]=\int_{\mathbb{R}}D(x,t)\nu_{m}(dx).$
-Then, by (2.4), $2G_{m}(t)\equiv F_{m}^{\prime}(t)$ and hence, $G_{m}(-t)>0>G_{m}(t)$ for every $t>0$.
-We show that
+$G_{m}(t)\coloneqq E\left[D(X_{1},t)\right]=\int_{\mathbb{R}}D(x,t)\nu_{m}(dx).$
+Then, by (2.6), $-2G_{m}(t)\equiv F_{m}^{\prime}(t)$ and hence, $G_{m}(-t)>0>G_{m}(t)$ for every $t>0$.
+We show the following:
 Lemma 3.1.
@@
 In particular,
-$\lim_{n\to\infty}\max_{t\in[-1,1]}\left|D_{n}(t)-G_{m}(t)\right|=0,\ \textup{ $P$-a.s.}$
+$\lim_{n\to\infty}\max_{t\in[-1,1]}\left|D_{n}(t)-G_{m}(t)\right|=0,\ \textup{ $\displaystyle P$-a.s.}$
 The constant $c_{\epsilon,1}$ is independent of $m$.
@@
 Let $t\in[-1,1]$.
-Let $Y_{i}:=D(X_{i},t)-G_{m}(t)$.
-Since $|D(X_{i},t)|\leq 1/2$ and $|G_{m}(t)|\leq 1/2$, $(Y_{i})_{i}$ are i.i.d., $|Y_{i}|\leq 1$ and $E[Y_{i}]=0$.
-By the Azuma-Hoeffding inequality (see Petrov [11, 2.6.2]),
+Let $Y_{i}\coloneqq D(X_{i},t)-G_{m}(t)$.
+Since $|D(X_{i},t)|\leq 1/2$ and $|G_{m}(t)|\leq 1/2$, it holds that $(Y_{i})_{i}$ are i.i.d., $|Y_{i}|\leq 1$ and $E[Y_{i}]=0$.
+By the Azuma-Hoeffding inequality (see Petrov [16, 2.6.2] or Boucheron, Lugosi and Massart [8, Theorem 2.8]),
 $P\left(\left|D_{n}(t)-G_{m}(t)\right|>\epsilon\right)=P\left(\left|\sum_{i=1}^{n}Y_{i}\right|>n\epsilon\right)\leq 2\exp\left(-\frac{\epsilon^{2}}{2}n\right).$
-Let $\mathcal{D}_{N}:=\{\ell/N:-N\leq\ell\leq N\}$ for $N\in\mathbb{N}$. Since $t\mapsto D(x,t)$ is Lipschitz continuous with the Lipschitz constant $1$,
+Let $\mathcal{D}_{N}\coloneqq\{\ell/N:-N\leq\ell\leq N\}$ for $N\in\mathbb{N}$. Since $t\mapsto D(x,t)$ is Lipschitz continuous with the Lipschitz constant $1$,
 $\max_{t\in[-1,1]}\left|D_{n}(t)-G_{m}(t)\right|\leq\max_{t\in\mathcal{D}_{N}}\left|D_{n}(t)-G_{m}(t)\right|+\frac{2}{N}.$
@@
 $G_{m}^{\prime}(t)=\int_{\mathbb{R}}\partial_{t}D(x,t)\nu_{m}(dx).$
-By the Lebesgue convergence theorem, $G_{m}^{\prime}$ is continuous. Hence, it suffices to show that $G_{m}^{\prime}(0)<0$.
+By the Lebesgue convergence theorem again, we see that $G_{m}^{\prime}$ is continuous. Hence, it suffices to show that $G_{m}^{\prime}(0)<0$.
 By the change of variables $x=\tan\theta$,
@@
 $\int_{0}^{\pi/4}\cos^{2m}\theta\cos(2\theta)d\theta-\int_{0}^{\pi/4}\cos^{2m}\left(\frac{\pi}{2}-\theta\right)\cos(2\theta)d\theta>0.$
-We also deal with the derivatives of $D_{n}(t)$ and $G_{m}(t)$ with respect to $t$. The following corresponds to [2, (3.32)].
-Lemma 3.3.
+We also deal with the derivatives of $D_{n}(t)$ and $G_{m}(t)$ with respect to $t$. The following corresponds to [3, (3.32)].
+Lemma $3.$
 For every $\epsilon>0$, there exists a positive constant $c_{\epsilon,2}$ depending on $\epsilon$ such that for every $n\geq 1$,
@@
 In particular,
-$\lim_{n\to\infty}\max_{t\in[-r_{m,2},r_{m,2}]}\left|D_{n}^{\prime}(t)-G_{m}^{\prime}(t)\right|=0,\ \textup{ $P$-a.s.}$
+$\lim_{n\to\infty}\max_{t\in[-r_{m,2},r_{m,2}]}\left|D_{n}^{\prime}(t)-G_{m}^{\prime}(t)\right|=0,\ \textup{ $\displaystyle P$-a.s.}$
 As in Lemma 3.3, the constant $c_{\epsilon,2}$ is also independent of $m$.
 Proof.
-By (3.7) below, $|\partial_{t}^{2}D(x,t)|\leq 3$, and hence, the map $t\mapsto\partial_{t}D(x,t)$ is Lipschitz continuous with the Lipschitz constant $3$.
-Let $Y^{\prime}_{i}:=\partial_{t}D(X_{i},t)-G_{m}^{\prime}(t)$.
+By
+(3.1) $\partial_{t}^{2}D(x,t)=\frac{2(x-t)((x-t)^{2}-3)}{(1+(x-t)^{2})^{3}},$
+$|\partial_{t}^{2}D(x,t)|\leq 3$, and hence, the map $t\mapsto\partial_{t}D(x,t)$ is Lipschitz continuous with the Lipschitz constant $3$.
+Let $Y^{\prime}_{i}\coloneqq\partial_{t}D(X_{i},t)-G_{m}^{\prime}(t)$.
 Since $|\partial_{t}D(X_{i},t)|\leq 1$ and $|G_{m}^{\prime}(t)|\leq 1$, $\left(Y^{\prime}_{i}\right)_{i}$ are i.i.d., $|Y_{i}^{\prime}|\leq 2$ and $E[Y_{i}^{\prime}]=0$.
@@
 Proposition 3.4.
-$P$-a.s. $\omega$, there exists $N_{5}(\omega)\in\mathbb{N}$ such that for every $n>N_{5}(\omega)$, $|C_{Q_{n}(\omega)}\cap[-r_{m,2},r_{m,2}]|=1$.
+$P$-a.s. $\omega$, there exists $N_{5}(\omega)\in\mathbb{N}$ such that for every $n>N_{5}(\omega)$, $\left|C_{Q_{n}(\omega)}\cap[-r_{m,2},r_{m,2}]\right|=1$.
 Proof.
@@
 Let
-$c_{m,2}:=\frac{1}{2}\min_{t\in[-r_{m,2},r_{m,2}]}-G_{m}^{\prime}(t),$
+$c_{m,2}\coloneqq\frac{1}{2}\min_{t\in[-r_{m,2},r_{m,2}]}-G_{m}^{\prime}(t),$
 which is positive by Lemma 3.2.
@@
 $\max_{t\in[-r_{m,2},r_{m,2}]}D_{n}^{\prime}(t)(\omega)\leq-c_{m,2},$
-in particular, $D_{n}(t)(\omega)$ is strictly decreasing in $t$ on $r_{m,2},r_{m,2}]$.
+in particular, $D_{n}(t)(\omega)$ is strictly decreasing in $t$ on $[-r_{m,2},r_{m,2}]$.
 Furthermore, by Lemma 3.1, it holds that $P$-a.s. $\omega$, there exists $N_{7}(\omega)\in\mathbb{N}$ such that for every $n>N_{7}(\omega)$,
@@
 Let
-$\mathcal{A}_{n}:=\mathcal{A}_{n,1}\cap\mathcal{A}_{n,2}\cap\mathcal{A}_{n,3}.$
-Let $\widetilde{\mathcal{A}_{i}}:=\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n,i},i=1,2,3$.
-By Lemma 3.1, $P\left(\widetilde{\mathcal{A}_{1}}\right)=1$.
-By Lemma 3.3, $P\left(\widetilde{\mathcal{A}_{2}}\right)=1$.
-By Propositions 2.7 and 3.4, $P\left(\widetilde{\mathcal{A}_{3}}\right)=1$.
-Since $\widetilde{\mathcal{A}_{1}}\cap\widetilde{\mathcal{A}_{2}}\cap\widetilde{\mathcal{A}_{3}}=\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}$, $P\left(\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}\right)=1$, and in particular, $\lim_{n\to\infty}P(\mathcal{A}_{n})=1$.
+(3.2) $\mathcal{A}_{n}\coloneqq\mathcal{A}_{n,1}\cap\mathcal{A}_{n,2}\cap\mathcal{A}_{n,3}.$
+Let
+$\widetilde{\mathcal{A}_{i}}\coloneqq\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n,i},\ \ i=1,2,3.$
+By Lemma 3.1, $\displaystyle P\left(\widetilde{\mathcal{A}_{1}}\right)=1$.
+By Lemma 3.3, $\displaystyle P\left(\widetilde{\mathcal{A}_{2}}\right)=1$.
+By Propositions 2.7 and 3.4, $\displaystyle P\left(\widetilde{\mathcal{A}_{3}}\right)=1$.
+Since $\displaystyle\widetilde{\mathcal{A}_{1}}\cap\widetilde{\mathcal{A}_{2}}\cap\widetilde{\mathcal{A}_{3}}=\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}$, $\displaystyle P\left(\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}\right)=1$, and in particular, $\displaystyle\lim_{n\to\infty}P(\mathcal{A}_{n})=1$.
 For every $t\in(-r_{m,2}/2,r_{m,2}/2)$, on $\mathcal{A}_{n}$, $\hat{\theta}_{n}<t$ if and only if $D_{n}(t)<0$.
-Let $y\in\mathbb{R}$. Then,
+Let $y\in\mathbb{R}$. Then
 $\lim_{n\to\infty}P\left(\sqrt{n}\hat{\theta}_{n}<y\right)-P\left(\left\{\sqrt{n}\hat{\theta}_{n}<y\right\}\cap\mathcal{A}_{n}\right)=0,$
@@
 Hence, it suffices to show that
-$\lim_{n\to\infty}P\left(D_{n}\left(\frac{y}{\sqrt{n}}\right)<0\right)=\int_{-\infty}^{y}\varphi_{m}(t)dt,$ (3.1)
-where $\varphi_{m}$ is the density function of the distribution $N\left(0,\frac{m+1}{m(2m-1)}\right)$.
+(3.3) $\lim_{n\to\infty}P\left(D_{n}\left(\frac{y}{\sqrt{n}}\right)<0\right)=\int_{-\infty}^{y}\varphi_{m}(t)dt,$
+where $\varphi_{m}$ is the density function of the distribution $\displaystyle N\left(0,\frac{m+1}{m(2m-1)}\right)$.
 It holds that
@@
 $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right).$
-By the symmetry,
-$E[D(X_{1},0)]=c_{m}\int_{\mathbb{R}}\frac{x}{(1+x^{2})^{1+m}}dx=0.$ (3.2)
+By symmetry,
+(3.4) $E\left[D(X_{1},0)\right]=c_{m}\int_{\mathbb{R}}\frac{x}{(1+x^{2})^{1+m}}dx=0.$
 By the change of variables $x=\tan\theta$,
-$E\left[D(X_{1},0)^{2}\right]=c_{m}\int_{\mathbb{R}}\frac{x^{2}}{(1+x^{2})^{2+m}}dx=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}=\frac{2m-1}{4m(m+1)},$ (3.3)
+(3.5) $E\left[D(X_{1},0)^{2}\right]=c_{m}\int_{\mathbb{R}}\frac{x^{2}}{(1+x^{2})^{2+m}}dx=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}=\frac{2m-1}{4m(m+1)},$
 where $B(\cdot,\cdot)$ is the beta function. Hence,
-$\sqrt{n}D_{n}\left(0\right)\Rightarrow N\left(0,\frac{2m-1}{4m(m+1)}\right),\ n\to\infty,$ (3.4)
+(3.6) $\sqrt{n}D_{n}\left(0\right)\Rightarrow N\left(0,\frac{2m-1}{4m(m+1)}\right),\ n\to\infty,$
 where $\Rightarrow$ denotes the convergence in distribution.
@@
 and
-$E\left[\partial_{t}D(X_{1},0)\right]$ $=c_{m}\int_{\mathbb{R}}\frac{x^{2}-1}{(1+x^{2})^{2+m}}dx$
-$=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}-\frac{B(1/2,m+3/2)}{B(1/2,m-1/2)}=-\frac{2m-1}{2(m+1)}.$ (3.5)
+$\displaystyle E\left[\partial_{t}D(X_{1},0)\right]$ $\displaystyle=c_{m}\int_{\mathbb{R}}\frac{x^{2}-1}{(1+x^{2})^{2+m}}dx$
+(3.7) $\displaystyle=\frac{B(3/2,m+1/2)}{B(1/2,m-1/2)}-\frac{B(1/2,m+3/2)}{B(1/2,m-1/2)}=-\frac{2m-1}{2(m+1)}.$
 Hence, by the strong law of large numbers,
-$\lim_{n\to\infty}\frac{y}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0)=-\frac{2m-1}{2(m+1)}y,\ \textup{ $P$-a.s.}$ (3.6)
-Since
-$\partial_{t}^{2}D(x,t)=\frac{2(x-t)((x-t)^{2}-3)}{(1+(x-t)^{2})^{3}},$ (3.7)
-$\max_{x,t\in\mathbb{R}}\left|\partial_{t}^{2}D(x,t)\right|=\max_{y\in\mathbb{R}}\frac{2|y||y^{2}-3|}{(1+y^{2})^{3}}=:C_{1}<\infty.$
+(3.8) $\lim_{n\to\infty}\frac{y}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0)=-\frac{2m-1}{2(m+1)}y,\ \textup{ $\displaystyle P$-a.s.}$
+By (3.1),
+$\max_{x,t\in\mathbb{R}}\left|\partial_{t}^{2}D(x,t)\right|=\max_{y\in\mathbb{R}}\frac{2|y||y^{2}-3|}{(1+y^{2})^{3}}\eqqcolon C_{1}<\infty.$
 By this and the mean value theorem, it holds [^3] This holds without any exceptional set. that
-$\left|D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right|\leq C_{1}\frac{|y|^{2}}{n}.$ (3.8)
+(3.9) $\left|D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right|\leq C_{1}\frac{|y|^{2}}{n}.$
 Hence,
-$\lim_{n\to\infty}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right)=0,\ \textup{ $P$-a.s.}$ (3.9)
-By (3.4), (3.6), (3.9) and Slutsky’s theorem,
+(3.10) $\lim_{n\to\infty}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(D\left(X_{i},\frac{y}{\sqrt{n}}\right)-D\left(X_{i},0\right)-\frac{y}{\sqrt{n}}\partial_{t}D(X_{i},0)\right)=0,\ \textup{ $\displaystyle P$-a.s.}$
+By (3.6), (3.8), (3.10) and Slutsky’s lemma,
 $\sqrt{n}D_{n}\left(\frac{y}{\sqrt{n}}\right)\Rightarrow N\left(-\frac{2m-1}{2(m+1)}y,\frac{2m-1}{4m(m+1)}\right),\ n\to\infty.$
-We see that (3.1) holds.
+Thus we see that (3.3) holds and the proof of Theorem 1.2 is completed.
+Subsection 8.2 below provides numerical verifications of Theorem 1.2 by using the Kolmogorov–Smirnov distance.
+Subsection 8.3 below provides confidence intervals of the parameter $\theta$ by using $\hat{\theta}_{n}$.
 Remark 3.5.
@@
 $\lim_{n\to\infty}P\left(\left|\{t\in\mathbb{R}|D_{n}(t)=0\}\right|=2k+1\right)=\exp\left(-c_{m}\right)\frac{c_{m}^{k}}{k!}.$
+(iii) We can apply [22, Theorem 5.23].
+We confirm the assumptions of the assertion.
+Let $\mathfrak{m}(x,\theta)\coloneqq-\log(1+(x-\theta)^{2}),\ x,\theta\in\mathbb{R}$.
+Then the MLE $\hat{\theta}_{n}$ maximizes the map $\theta\mapsto\sum_{i=1}^{n}\mathfrak{m}(X_{i},\theta)$.
+We see that $\mathfrak{m}\in C^{\infty}(\mathbb{R}^{2})$ and by (2.3), $|\mathfrak{m}(x,\theta_{1})-\mathfrak{m}(x,\theta_{2})|\leq|\theta_{1}-\theta_{2}|$.
+It holds that $E\left[\mathfrak{m}(X_{1},\theta)\right]=-F_{m}(\theta)$.
+Since $F_{m}^{\prime}=-2G_{m}$ and the argument in the proof of Lemma 3.2, $G_{m}\in C^{1}(\mathbb{R})$ and hence $F_{m}\in C^{2}(\mathbb{R})$.
+Hence $F_{m}$ admits the second-order Taylor expansion at $\theta=0$.
+Furthermore $F_{m}^{\prime\prime}(0)=-2G_{m}^{\prime}(0)>0$ by Lemma 3.2.
+Finally we recall Theorem $1.$ Thus the assumptions of [22, Theorem 5.23] are satisfied and Theorem 1.2 follows.
 4. Proof of Theorem 1.3
-We follow the strategy of [3, Section 5]. As in the case of Theorem 1.2, we cannot apply the result of [3] directly and need to modify several parts.
-Let $\phi(n):=\sqrt{2n\log\log n}$. Let
-$R_{n}:=\frac{1}{\phi(n)}\sum_{i=1}^{n}D(X_{i},0),$
-$S_{n}:=\frac{1}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0),$
+We first give an outline of the proof.
+We follow the strategy of Barczy and Páles [4, Section 5].
+As in the case of Theorem 1.2, we cannot apply the result of [4] directly and need to modify several parts.
+We evaluate the normalized score along the scale of the law of the iterated logarithm and use the decomposition in (4.1) below into the leading fluctuation term $R_{n}$, the drift term $\gamma S_{n}$, and the remainder $T_{n}^{(\gamma)}$.
+We apply the Kolmogorov law of the iterated logarithm to $R_{n}$, the strong law of large numbers to $S_{n}$, and finally show that $T_{n}^{(\gamma)}$ is negligible.
+Let $\phi(n)\coloneqq\sqrt{2n\log\log n}$. Let
+$R_{n}\coloneqq\frac{1}{\phi(n)}\sum_{i=1}^{n}D(X_{i},0),$
+$S_{n}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\partial_{t}D(X_{i},0),$
 and
-$T_{n}^{(\gamma)}:=\frac{1}{\phi(n)}\sum_{i=1}^{n}\left(D\left(X_{i},\gamma\frac{\phi(n)}{n}\right)-D(x_{i},0)-\gamma\frac{\phi(n)}{n}\partial_{t}D(X_{i},0)\right).$
-Then,
-$R_{n}+\gamma S_{n}+T_{n}^{(\gamma)}=\frac{1}{\phi(n)}\sum_{i=1}^{n}D\left(X_{i},\gamma\frac{\phi(n)}{n}\right).$ (4.1)
-Since $\max_{x,t\in\mathbb{R}}|D(x,t)|\leq 1/2$, by the Kolmogorov law of the iterated logarithm, there exists an event $\Omega_{4}$ such that $P(\Omega_{4})=1$ and for every $\omega\in\Omega_{4}$,
-$\limsup_{n\to\infty}R_{n}(\omega)=\sqrt{\frac{2m-1}{4m(m+1)}}.$ (4.2)
+$T_{n}^{(\gamma)}\coloneqq\frac{1}{\phi(n)}\sum_{i=1}^{n}\left(D\left(X_{i},\gamma\frac{\phi(n)}{n}\right)-D(X_{i},0)-\gamma\frac{\phi(n)}{n}\partial_{t}D(X_{i},0)\right).$
+Then
+(4.1) $R_{n}+\gamma S_{n}+T_{n}^{(\gamma)}=\frac{1}{\phi(n)}\sum_{i=1}^{n}D\left(X_{i},\gamma\frac{\phi(n)}{n}\right).$
+Since $\displaystyle\max_{x,t\in\mathbb{R}}|D(x,t)|\leq 1/2$, by the Kolmogorov law of the iterated logarithm, there exists an event $\Omega_{4}$ such that $P(\Omega_{4})=1$ and for every $\omega\in\Omega_{4}$,
+(4.2) $\limsup_{n\to\infty}R_{n}(\omega)=\sqrt{\frac{2m-1}{4m(m+1)}}.$
 By the strong law of large numbers, there exists an event $\Omega_{5}$ such that $P(\Omega_{5})=1$ and for every $\omega\in\Omega_{5}$,
-$\lim_{n\to\infty}S_{n}(\omega)=-\frac{2m-1}{2(m+1)}.$ (4.3)
-Let $\Omega_{6}:=\Omega_{4}\cap\Omega_{5}\cap\left(\cup_{N\geq 1}\cap_{n\geq N}\mathcal{A}_{n}\right)$. Then, $P\left(\Omega_{6}\right)=1$.
+(4.3) $\lim_{n\to\infty}S_{n}(\omega)=-\frac{2m-1}{2(m+1)}.$
+Let $\displaystyle\Omega_{6}\coloneqq\Omega_{4}\cap\Omega_{5}\cap\left(\bigcup_{N\geq 1}\bigcap_{n\geq N}\mathcal{A}_{n}\right)$. Then $P\left(\Omega_{6}\right)=1$.
 By the uniform estimate (3.9), for every $\omega\in\Omega_{6}$ and every $\gamma\in\mathbb{R}$,
-$\lim_{n\to\infty}T_{n}^{(\gamma)}(\omega)=0.$ (4.4)
-For notational convenience, let $\sigma_{m}:=\sqrt{\dfrac{m+1}{m(2m-1)}}$.
+(4.4) $\lim_{n\to\infty}T_{n}^{(\gamma)}(\omega)=0.$
+Let $\sigma_{m}\coloneqq\sqrt{\dfrac{m+1}{m(2m-1)}}$.
 Let $\omega\in\Omega_{6}$ and $\epsilon>0$.
-Then, there exists $N_{9}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{9}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)<(\sigma_{m}+\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}+\epsilon)\frac{\phi(n)}{n}\right)<0$ holds.
+Then there exists $N_{9}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{9}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)<(\sigma_{m}+\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\displaystyle\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}+\epsilon)\frac{\phi(n)}{n}\right)<0$ holds.
 By (4.1), this is equivalent to
-$R_{n}(\omega)+(\sigma_{m}+\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}+\epsilon)}(\omega)<0.$ (4.5)
+(4.5) $R_{n}(\omega)+(\sigma_{m}+\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}+\epsilon)}(\omega)<0.$
 By (4.2), (4.3) and (4.4), there exists $N_{10}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{10}(\omega,\epsilon)$, (4.5) holds.
@@
 By letting $\epsilon\to 0$,
-$\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\leq\sigma_{m}.$ (4.6)
+(4.6) $\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\leq\sigma_{m}.$
 We can show the lower bound in the same manner.
-There exists $N_{11}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{11}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)>(\sigma_{m}-\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}-\epsilon)\frac{\phi(n)}{n}\right)>0$ holds.
+There exists $N_{11}(\omega,\epsilon)\in\mathbb{N}$ such that for every $n\geq N_{11}(\omega,\epsilon)$, $\hat{\theta}_{n}(\omega)>(\sigma_{m}-\epsilon)\dfrac{\phi(n)}{n}$ holds if and only if $\displaystyle\sum_{i=1}^{n}D\left(X_{i}(\omega),(\sigma_{m}-\epsilon)\frac{\phi(n)}{n}\right)>0$ holds.
 By (4.1), this is equivalent to
-$R_{n}(\omega)+(\sigma_{m}-\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}-\epsilon)}(\omega)>0.$ (4.7)
+(4.7) $R_{n}(\omega)+(\sigma_{m}-\epsilon)S_{n}(\omega)+T_{n}^{(\sigma_{m}-\epsilon)}(\omega)>0.$
 By (4.2), (4.3) and (4.4), (4.7) holds for infinitely many $n$. Hence, $\hat{\theta}_{n}(\omega)>(\sigma_{m}-\epsilon)\dfrac{\phi(n)}{n}$ holds for infinitely many $n$, and hence,
@@
 By letting $\epsilon\to 0$,
-$\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\geq\sigma_{m}.$ (4.8)
+(4.8) $\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}\geq\sigma_{m}.$
 By (4.6) and (4.8),
 $\limsup_{n\to\infty}\frac{n\hat{\theta}_{n}(\omega)}{\phi(n)}=\sigma_{m}.$
-This completes the proof.
+This completes the proof of Theorem 1.3.
 5. Proof of Theorem 1.4
-We prove Theorem 1.4 by following the strategy of [2].
-In the above section, we have seen that for $i=1,2$, $P(\mathcal{A}_{n,i}^{c})$ decays exponentially fast.
-Here we show that $P\left(\mathcal{A}_{n,3}^{c}\right)$ also decays exponentially fast.
-Recall the definition of $F_{m}$ in (2.1). For notational convenience, let $\widetilde{F}_{m}(t):=\exp\left(F_{m}(t)\right)$.
+We first give an outline of the proof.
+We prove Theorem 1.4 by following the strategy of Bai and Fu [3].
+We first recall that $\left|P\left(\hat{\theta}_{n}>\epsilon\right)-P(D_{n}(\epsilon)>0)\right|\leq 2P\left(\mathcal{A}_{n}^{c}\right)$ by the arguments of Section 3.
+The first step is to show (5.6) below, which states that the probability of the localization event $P\left(\mathcal{A}_{n}^{c}\right)$ decays exponentially fast as $n\to\infty$.
+In Section 3, we have seen that for $i=1,2$, $P(\mathcal{A}_{n,i}^{c})$ decays exponentially fast.
+So it remains to control $P\left(\mathcal{A}_{n,3}^{c}\right)$.
+This is done by two large deviation estimates for $L_{n}(t)$.
+First, Lemma 5.3 below controls $\displaystyle\inf_{|t|\geq r}L_{n}(t)$ by using Lemmas 5.1 and 5.2 below.
+Second, Lemma 5.4 below controls $L_{n}(0)$ by an exponential Chebyshev bound.
+Combining these bounds, we obtain (5.6).
+After this, the problem reduces to estimating $P(D_{n}(\epsilon)>0)$.
+We center $D(X_{i},\epsilon)$ by its mean $G_{m}(\epsilon)$ and then apply two deviation inequalities in Lemmas 5.6 and 5.7 below with its variance $H_{m}(\epsilon)$.
+The Taylor expansions of $G_{m}$ and $H_{m}$ around $0$ in Lemma 5.5 below together with (5.10) identify the quadratic rate constant, which matches the upper and lower bounds in (1.1) and (1.2) respectively.
+Recall the definition of $F_{m}$ in (2.1). Let
+$\widetilde{F}_{m}(t)\coloneqq\exp\left(F_{m}(t)\right).$
 Lemma 5.1.
@@
 The statement is equivalent to
-$\lim_{t\to\infty}F_{m}(t)-\log(1+t^{2})=0.$ (5.1)
+(5.1) $\lim_{t\to\infty}F_{m}(t)-\log(1+t^{2})=0.$
 We see that
@@
 Now we can apply the Lebesgue convergence theorem. ∎
-The following corresponds to [2, (3.15)].
+Let
+(5.2) $\lambda_{m}\coloneqq\frac{1}{2}\left(m-\frac{1}{2}\right)$
+and
+(5.3) $\delta_{m}(r)\coloneqq\frac{F_{m}(r)-F_{m}(0)}{4},\ r>0.$
+These definitions will be used frequently not only in this section but also in the following section.
+The following corresponds to [3, (3.15)].
 Lemma 5.2.
-Let $r>0$. Assume that
-$0<\delta<\min\left\{\frac{1}{2}\left(m-\frac{1}{2}\right),\frac{F_{m}(r)-F_{m}(0)}{4}\right\}.$
-Then, there exists a positive constant $c_{m,3}$ depending only on $m$ such that for every $t$ with $|t|>r$ and every $n\geq 1$,
+Let $r>0$.
+Assume that $0<\delta<\min\left\{\lambda_{m},\delta_{m}(r)\right\}$.
+Then there exists a positive constant $c_{m,3}$ depending only on $m$ such that for every $t$ with $|t|>r$ and every $n\geq 1$,
 $P(L_{n}(t)\leq F_{m}(0)+\delta)\leq\exp\left(-\frac{\delta}{c_{m,3}}(F_{m}(t)-F_{m}(0)-2\delta)n\right).$
+The following proof is similar to the proof of Bernstein’s inequality ([8, Theorem 2.10]).
+However the estimates are different, see (5.4) below.
+The proof below is easier in the sense that there is no need to consider the Fenchel–Legendre transform.
 Proof.
-We assume that $t>r$. The proof is the same for the case that t<rt<-r. We see that
+We assume that $t>r$. The proof is the same for the case that $t<-r$. We see that
 $P(L_{n}(t)\leq F_{m}(0)+\delta)=P\left(\sum_{i=1}^{n}(F_{m}(t)-\log(1+(X_{i}-t)^{2}))\geq n(F_{m}(t)-F_{m}(0)-\delta)\right).$
@@
 for every $\lambda>0$.
-Assume that $0<\lambda<m-1/2$. Then,
-$E\left[\exp\left(\lambda\left|F_{m}(t)-\log(1+(X_{1}-t)^{2})\right|\right)\right]\leq\exp(\lambda F_{m}(t))E\left[(1+(X_{1}-t)^{2})^{\lambda}\right]<\infty.$
+Assume that $0<\lambda<m-\dfrac{1}{2}$. Then
+$E\left[\exp\left(\lambda\left|F_{m}(t)-\log(1+(X_{1}-t)^{2})\right|\right)\right]\leq\exp(\lambda F_{m}(t))E\left[\left(1+(X_{1}-t)^{2}\right)^{\lambda}\right]<\infty.$
 Therefore, we can apply the Taylor expansion and obtain that
-$E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]=\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right].$
-Since $E\left[\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right]=0$,
-$\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]=1+\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right].$
+$E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]=\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right].$
+where we let $\displaystyle\Psi(x,t)\coloneqq\log\frac{\widetilde{F}_{m}(t)}{1+(x-t)^{2}}$. Since $E\left[\Psi(X_{1},t)\right]=0$,
+$\sum_{k=0}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right]=1+\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right].$
 By Lemma 5.1,
-$c_{m,4}:=\sup_{x\leq t/2,t\geq 0}\log\frac{\widetilde{F}_{m}(t)}{1+(x-t)^{2}}<\infty.$
+$c_{m,4}\coloneqq\sup_{x\leq t/2,t\geq 0}\Psi(x,t)<\infty.$
 Hence,
-$E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]\leq c_{m,4}^{k}P(X_{1}\leq t/2)+F_{m}(t)^{k}P(X_{1}>t/2).$
+(5.4) $E\left[\Psi(X_{1},t)^{k}\right]\leq c_{m,4}^{k}P(X_{1}\leq t/2)+F_{m}(t)^{k}P(X_{1}>t/2).$
 Since
 $P(X_{1}>t/2)\leq c_{m}\int_{t/2}^{\infty}x^{-2m}dx\leq c_{m}4^{m}t^{1-2m},$
-$E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]\leq c_{m,4}^{k}+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{k},$
-where we let $c_{m,5}:=c_{m}4^{m}$. Hence,
-$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]$
-$\leq\frac{\lambda^{2}}{2}\left(c_{m,4}^{2}\exp(\lambda c_{m,4})+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda F_{m}(t))\right).$
-Since $0<\lambda<m-\frac{1}{2}$,
+we see that
+$E\left[\Psi(X_{1},t)^{k}\right]\leq c_{m,4}^{k}+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{k},$
+where we let $c_{m,5}\coloneqq c_{m}4^{m}$. Hence,
+$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right]\leq\frac{\lambda^{2}}{2}\left(c_{m,4}^{2}\exp(\lambda c_{m,4})+\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda F_{m}(t))\right).$
+Since $0<\lambda<m-\dfrac{1}{2}$,
 $\lim_{t\to\infty}\frac{F_{m}(t)^{2}\exp(\lambda F_{m}(t))}{t^{2m-1}}=0,$
@@
 $\sup_{t\geq 0}\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda F_{m}(t))<\infty.$
-Let $\lambda_{m}:=\frac{1}{2}(m-\frac{1}{2})$. Then, for every $\lambda\in(0,\lambda_{m})$,
-$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\left(\log\frac{\widetilde{F}_{m}(t)}{1+(X_{1}-t)^{2}}\right)^{k}\right]\leq\lambda^{2}c_{m,6},$
+Recall (5.2). Then, for every $\lambda\in(0,\lambda_{m})$,
+$\sum_{k=2}^{\infty}\frac{\lambda^{k}}{k!}E\left[\Psi(X_{1},t)^{k}\right]\leq\lambda^{2}c_{m,6},$
 where we let
-$c_{m,6}:=\frac{1}{2}\left(c_{m,4}^{2}\exp(\lambda_{m}c_{m,4})+\sup_{t\geq 0}\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda_{m}F_{m}(t))\right)<\infty.$
+$c_{m,6}\coloneqq\frac{1}{2}\left(c_{m,4}^{2}\exp(\lambda_{m}c_{m,4})+\sup_{t\geq 0}\min\left\{1,\frac{c_{m,5}}{t^{2m-1}}\right\}F_{m}(t)^{2}\exp(\lambda_{m}F_{m}(t))\right)<\infty.$
 We can assume that $c_{m,6}\geq 1$ because if $c_{m,6}<1$, then we can replace $c_{m,6}$ with $c_{m,6}+1$.
-Therefore, for every $\lambda\in\left(0,\min\left\{\lambda_{m},\frac{F_{m}(r)-F_{m}(0)}{4}\right\}\right)$,
+Therefore, for every $\displaystyle\lambda\in\left(0,\min\left\{\lambda_{m},\delta_{m}(r)\right\}\right)$,
 $E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]\leq\exp(\lambda^{2}c_{m,6}).$
-If we let $\lambda:=\delta/c_{m,6}$, then, $0<\lambda<m-1/2$, and,
+If we let $\lambda\coloneqq\delta/c_{m,6}$, then, $0<\lambda<m-\dfrac{1}{2}$, and,
 $\exp\left(-\lambda(F_{m}(t)-F_{m}(0)-\delta)\right)E\left[\exp\left(\lambda(F_{m}(t)-\log(1+(X_{1}-t)^{2}))\right)\right]$
@@
 Thus, the assertion holds for $c_{m,3}=c_{m,6}$. ∎
-Let $\lambda_{m}(r):=\frac{1}{2}\min\left\{\frac{1}{2}\left(m-\frac{1}{2}\right),\frac{F_{m}(r)-F_{m}(0)}{4}\right\}$.
-The following corresponds to [2, (3.21)] [^4] There is a typo in [2, (3.21)]. The supremum in [2, (3.21)] should be the infimum..
+Let $\displaystyle\lambda_{m}(r)\coloneqq\frac{1}{2}\min\left\{\lambda_{m},\delta_{m}(r)\right\}$.
+The following corresponds to [3, (3.21)] [^4] There is a typo in [3, (3.21)]. The supremum in [3, (3.21)] should be the infimum..
 Lemma 5.3.
-Let $r>0$. Assume that $0<\delta<\lambda_{m}(r)$. Then, there exists $N(r,\delta)\in\mathbb{N}$ such that for every $n\geq N(r,\delta)$,
+Let $r>0$. Assume that $\displaystyle 0<\delta<\lambda_{m}(r)$. Then there exists $N(r,\delta)\in\mathbb{N}$ such that for every $n\geq N(r,\delta)$,
 $P\left(\inf_{|t|\geq r}L_{n}(t)<F_{m}(0)+\delta\right)\leq 2\exp\left(-\frac{\delta^{2}}{8c_{m,3}}n\right),$
 where $c_{m,3}$ is the constant appearing in Lemma 5.2.
-We remark that $r>0$ can be taken arbitrarily small.
+We remark that $r>0$ can be taken arbitrarily small. We first discretize $[r,\infty)$ by the Lipschitz continuity of $L_{n}(t)$ and then apply Lemma 5.2.
 Proof.
 We show that
-$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)\leq\exp\left(-\frac{\delta^{2}}{8c_{m,3}}n\right).$ (5.2)
-Since $L_{n}^{\prime}(t)=-2D_{n}(t)$ and $|D_{n}(t)|\leq 1/2$,
+(5.5) $P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)\leq\exp\left(-\frac{\delta^{2}}{8c_{m,3}}n\right).$
+Since $L_{n}^{\prime}(t)=-2D_{n}(t)$ and $|D_{n}(t)|\leq 1/2$, it holds that
 $\left\{\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right\}\subset\bigcup_{k\geq 1}\left\{L_{n}(k\delta+r)<F_{m}(0)+2\delta\right\}$
@@
 The case that $t\leq-r$ can be dealt with in the same manner. ∎
-The following corresponds to [2, (3.25)] [^5] There is also a typo in [2, (3.25)].
-$n^{2}$” in the right hand side of the inequality in [2, (3.25)] should be $n\delta^{2}$”..
-Recall that $\lambda_{m}=\frac{1}{2}(m-\frac{1}{2})$.
+The following corresponds to [3, (3.25)] [^5] There is also a typo in [3, (3.25)].
+$n^{2}$” in the right hand side of the inequality in [3, (3.25)] should be $n\delta^{2}$”..
+Recall the definition of $\lambda_{m}$ in (5.2).
 Lemma 5.4.
@@
 $P(L_{n}(0)\geq F_{m}(0)+\delta)\leq\left(\exp(-\lambda\delta)E\left[\exp(\lambda(\log(1+X_{1}^{2})-F_{m}(0)))\right]\right)^{n}.$
-Since $E[\log(1+X_{1}^{2})]=F_{m}(0)$,
+Since $E\left[\log(1+X_{1}^{2})\right]=F_{m}(0)$,
 $E\left[\exp(\lambda(\log(1+X_{1}^{2})-F_{m}(0)))\right]\leq\exp\left(\frac{\lambda^{2}}{2}c_{m,7}\right),$
 where we let
-$c_{m,7}:=E\left[(\log(1+X_{1}^{2})-F_{m}(0)))^{2}\exp\left(\lambda_{m}\left|\log(1+X_{1}^{2})-F_{m}(0))\right|\right)\right].$
-Now let $\lambda:=\delta/c_{m,7}$. ∎
+$c_{m,7}\coloneqq E\left[(\log(1+X_{1}^{2})-F_{m}(0)))^{2}\exp\left(\lambda_{m}\left|\log(1+X_{1}^{2})-F_{m}(0))\right|\right)\right].$
+Now let $\lambda\coloneqq\delta/c_{m,7}$. ∎
 Let
-$c_{m,8}:=\dfrac{1}{2}\min\left\{\lambda_{m}(r_{m,2}/3),c_{m,7}\lambda_{m}\right\}.$
-Let $\mathcal{B}_{n,1}$ be the event that $\inf_{|t|\geq r_{m,2}/3}L_{n}(t)<F_{m}(0)+c_{m,8}$.
+$c_{m,8}\coloneqq\dfrac{1}{2}\min\left\{\lambda_{m}(r_{m,2}/3),c_{m,7}\lambda_{m}\right\}.$
+Let $\mathcal{B}_{n,1}$ be the event that $\displaystyle\inf_{|t|\geq r_{m,2}/3}L_{n}(t)<F_{m}(0)+c_{m,8}$.
 Let $\mathcal{B}_{n,2}$ be the event that $L_{n}(0)\geq F_{m}(0)+c_{m,8}$.
-Then, $\hat{\theta}_{n}\in[-r_{m,2}/2,r_{m,2}/2]$ on the event $\mathcal{B}_{n,1}\cap\mathcal{B}_{n,2}$.
+Then $\hat{\theta}_{n}\in[-r_{m,2}/2,r_{m,2}/2]$ on the event $\mathcal{B}_{n,1}\cap\mathcal{B}_{n,2}$.
 Therefore,
@@
 and hence,
-$\left|P\left(\hat{\theta}_{n}>\epsilon\right)-P(D_{n}(\epsilon)>0)\right|\leq 2P(\mathcal{A}_{n}^{c})\leq 2c_{m,9}\exp(-c_{m,10}n),\ n\geq 1.$ (5.3)
+(5.6) $\left|P\left(\hat{\theta}_{n}>\epsilon\right)-P(D_{n}(\epsilon)>0)\right|\leq 2P(\mathcal{A}_{n}^{c})\leq 2c_{m,9}\exp(-c_{m,10}n),\ n\geq 1.$
 Let
-$H_{m}(\epsilon):=\textup{Var}(D(X_{1},\epsilon))=E\left[D(X_{1},\epsilon)^{2}\right]-G_{m}(\epsilon)^{2}.$
+$H_{m}(\epsilon)\coloneqq\textup{Var}(D(X_{1},\epsilon))=E\left[D(X_{1},\epsilon)^{2}\right]-G_{m}(\epsilon)^{2}.$
 Lemma $5.$
-(1) $G_{m}(\epsilon)=-\frac{2m-1}{2(m+1)}\epsilon+O(\epsilon^{2}),\ \epsilon\to+0$.
-(2) $H_{m}(\epsilon)=\frac{2m-1}{4m(m+1)}+O(\epsilon),\ \epsilon\to+0$.
+It holds that
+(1) $\displaystyle G_{m}(\epsilon)=-\frac{2m-1}{2(m+1)}\epsilon+O(\epsilon^{2}),\ \epsilon\to+0$.
+(2) $\displaystyle H_{m}(\epsilon)=\frac{2m-1}{4m(m+1)}+O(\epsilon),\ \epsilon\to+0$.
 Proof.
 (1) By (3.9),
-$\left|D(X_{1},\epsilon)-D(X_{1},0)-\epsilon\partial_{t}D(X_{1},0)\right|\leq C_{1}\epsilon^{2}.$ (5.4)
+(5.7) $\left|D(X_{1},\epsilon)-D(X_{1},0)-\epsilon\partial_{t}D(X_{1},0)\right|\leq C_{1}\epsilon^{2}.$
 By (3.4) and (3),
-$E[D(X_{1},0)]=0,E\left[\partial_{t}D(X_{1},0)\right]=-\frac{2m-1}{2(m+1)}.$
+$E\left[D(X_{1},0)\right]=0,E\left[\partial_{t}D(X_{1},0)\right]=-\frac{2m-1}{2(m+1)}.$
 The estimate follows from these equalities and (5.7).
 (2) By (5.7), there exists a positive constant $C_{2}$ such that for every $\epsilon\in(0,1)$,
-$\left|D(X_{1},\epsilon)^{2}-D(X_{1},0)^{2}-2\epsilon D(X_{1},0)\partial_{t}D(X_{1},0)\right|\leq C_{2}\epsilon^{2}.$ (5.5)
+(5.8) $\left|D(X_{1},\epsilon)^{2}-D(X_{1},0)^{2}-2\epsilon D(X_{1},0)\partial_{t}D(X_{1},0)\right|\leq C_{2}\epsilon^{2}.$
 Since $D(X_{1},0)$ and $\partial_{t}D(X_{1},0)$ are bounded, $D(X_{1},0)\partial_{t}D(X_{1},0)$ is also bounded, and in particular, is integrable. By (3.5),
-$H_{m}(0)=E[D(X_{1},0)^{2}]=\frac{2m-1}{4m(m+1)}.$
+$H_{m}(0)=E\left[D(X_{1},0)^{2}\right]=\frac{2m-1}{4m(m+1)}.$
 The estimate follows from this equality and (5.8). ∎
-We show (i). We consider the asymptotics of $P(D_{n}(\epsilon)>0)$.
+We show (i) of Theorem 1.4. We consider the asymptotics of $P(D_{n}(\epsilon)>0)$ as $n\to\infty$.
 We first give the upper estimate. We remark that $|D(X_{i},\epsilon)-G_{m}(\epsilon)|\leq\frac{1}{2}-G_{m}(\epsilon)$ and by Lemma 5.5,
@@
 $\left|D(X_{i},\epsilon)-G_{m}(\epsilon)\right|\leq H_{m}(\epsilon).$
-Lemma 5.6 (Petrov [11, Lemma 7.1] [^6] The statement is a little different from [2, Lemma 1]. In [2, Lemma 1], this assertion holds for large $n$, but this is valid for every $n\geq 1$.).
-Let $Z_{i},i\geq 1$, be i.i.d. random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}:=\textup{Var}(Z_{1})>0$. Then, for every $n\geq 1$ and every $x\in[0,\sigma^{2}/M]$,
+Lemma 5.6 (Petrov [16, Lemma 7.1] [^6] The statement is a little different from [3, Lemma 1]. In [3, Lemma 1], this assertion holds for large $n$, but this is valid for every $n\geq 1$.).
+Let $Z_{i},i\geq 1$, be i.i.d.
+random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}\coloneqq\textup{Var}(Z_{1})>0$.
+Then, for every $n\geq 1$ and every $x\in[0,\sigma^{2}/M]$,
 $P\left(\sum_{i=1}^{n}Z_{i}\geq nx\right)\leq\exp\left(-\frac{nx^{2}}{2\sigma^{2}}\left(1-\frac{Mx}{2\sigma^{2}}\right)\right).$
 By this lemma, it holds that for every $\epsilon\in(0,\epsilon_{m,1})$ and every $n\geq 1$,
-$P(D_{n}(\epsilon)>0)$ $=P\left(\sum_{i=1}^{n}D(X_{i},\epsilon)-G_{m}(\epsilon)>-nG_{m}(\epsilon)\right)$
-$\leq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\right).$ (5.6)
+$\displaystyle P(D_{n}(\epsilon)>0)$ $\displaystyle=P\left(\sum_{i=1}^{n}D(X_{i},\epsilon)-G_{m}(\epsilon)>-nG_{m}(\epsilon)\right)$
+(5.9) $\displaystyle\leq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\right).$
 By Lemma 5.5,
-$\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\sim\frac{m(2m-1)}{m+1}\epsilon^{2},\ \epsilon\to+0,$ (5.7)
+(5.10) $\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\sim\frac{m(2m-1)}{m+1}\epsilon^{2},\ \epsilon\to+0,$
 in particular,
@@
 $\limsup_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon\right)}{n}\leq-\frac{G_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right).$
-By this, Lemma 5.5,
+By this, Lemma 5.5, and (5.10),
 $\limsup_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\limsup_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon\right)}{n}\right)\leq-\frac{m(2m-1)}{2(m+1)}.$
 The same argument is applicable to $P\left(\hat{\theta}_{n}<-\epsilon\right)$ and we obtain (1.1).
-We next give the lower estimate. By Lemma 5.5,
+We next give the lower estimate (1.2). By Lemma 5.5,
 $\lim_{\epsilon\to+0}G_{m}(\epsilon)=0\textup{ and }\lim_{\epsilon\to+0}H_{m}(\epsilon)=E\left[D(X_{1},0)^{2}\right]>0.$
-Lemma 5.7 (Petrov [11, Lemma 7.2] [^7] The statement is a little different from [11, Lemma 7.2], however, we can show this assertion in the same manner as in the proof of [11, Lemma 7.2].).
+Lemma 5.7 (Petrov [16, Lemma 7.2] [^7] The statement is a little different from [16, Lemma 7.2], however, we can show this assertion in the same manner as in the proof of [16, Lemma 7.2].).
 Let $Z_{i},i\geq 1$, be i.i.d.
-random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}:=\textup{Var}(Z_{1})>0$.
+random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}\coloneqq\textup{Var}(Z_{1})>0$.
 Then, for every $\eta>0$, there exists $r>0$ such that for every $x\in[0,r]$, there exists $N$ such that for every $n\geq N$,
@@
 By this lemma, for every $\eta>0$, there exists $\epsilon_{\eta}>0$ depending on $m$ and $\eta$ such that for every $\epsilon\in(0,\epsilon_{\eta})$, there exists $N_{\eta,\epsilon,1}\in\mathbb{N}$ such that for every $n\geq N_{\eta,\epsilon,1}$,
-$P(D_{n}(\epsilon)>0)\geq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}(1+\eta)\right).$ (5.8)
+(5.11) $P(D_{n}(\epsilon)>0)\geq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}(1+\eta)\right).$
 In the same manner as in the upper bound, it holds that there exists $\epsilon_{\eta,2}>0$ depending on $\eta$ such that for every $\epsilon\in(0,\epsilon_{\eta,2})$, there exists $N_{\eta,\epsilon,2}\in\mathbb{N}$ such that for every $n\geq N_{\eta,\epsilon,2}$,
@@
 $\liminf_{\epsilon\to+0}\frac{1}{\epsilon^{2}}\left(\liminf_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon\right)}{n}\right)\geq-\frac{m(2m-1)}{2(m+1)}.$
-The same argument is applicable to $P\left(\hat{\theta}_{n}<-\epsilon\right)$ and we obtain (1.2).
-Now we show (ii), but the proof is almost identical to the proof of (i).
+The same argument is applicable to $P\left(\hat{\theta}_{n}<-\epsilon\right)$ and we obtain (1.2). Thus the proof of (i) of Theorem 1.4 is completed.
+Now we show (ii) of Theorem 1.4, but the proof is almost identical to the proof of (i).
 By (5), it holds that for large $n$,
-$P(D_{n}(\epsilon/\lambda_{n})>0)\leq\exp\left(-\frac{nG_{m}(\epsilon/\lambda_{n})^{2}}{2H_{m}(\epsilon/\lambda_{n})}\left(1+\frac{G_{m}(\epsilon/\lambda_{n})}{2H_{m}(\epsilon/\lambda_{n})}\right)\right).$
+$P(D_{n}(\epsilon/a_{n})>0)\leq\exp\left(-\frac{nG_{m}(\epsilon/a_{n})^{2}}{2H_{m}(\epsilon/a_{n})}\left(1+\frac{G_{m}(\epsilon/a_{n})}{2H_{m}(\epsilon/a_{n})}\right)\right).$
 By Lemma 5.5,
-$\lim_{n\to\infty}\lambda_{n}^{2}\frac{G_{m}(\epsilon/\lambda_{n})^{2}}{H_{m}(\epsilon/\lambda_{n})}\left(1+\frac{G_{m}(\epsilon/\lambda_{n})}{2H_{m}(\epsilon/\lambda_{n})}\right)=\frac{m(2m-1)}{m+1}.$
+$\lim_{n\to\infty}a_{n}^{2}\frac{G_{m}(\epsilon/a_{n})^{2}}{H_{m}(\epsilon/a_{n})}\left(1+\frac{G_{m}(\epsilon/a_{n})}{2H_{m}(\epsilon/a_{n})}\right)=\frac{m(2m-1)}{m+1}.$
 Therefore, we obtain that
-$\limsup_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/\lambda_{n})>0\right)}{n/\lambda_{n}^{2}}\leq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$ (5.9)
+(5.12) $\limsup_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/a_{n})>0\right)}{n/a_{n}^{2}}\leq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
 By (5.11) and Lemma 5.5, we obtain that
-$\liminf_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/\lambda_{n})>0\right)}{n/\lambda_{n}^{2}}\geq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$ (5.10)
+(5.13) $\liminf_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/a_{n})>0\right)}{n/a_{n}^{2}}\geq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
 (5.12) and (5.13) imply that
-$\lim_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/\lambda_{n})>0\right)}{n/\lambda_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
+$\lim_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/a_{n})>0\right)}{n/a_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
 By this and (5.6),
-$\lim_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon/\lambda_{n}\right)}{n/\lambda_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
-$P\left(\hat{\theta}_{n}<-\epsilon/\lambda_{n}\right)$ can be dealt with in the same manner.
+$\lim_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon/a_{n}\right)}{n/a_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
+$P\left(\hat{\theta}_{n}<-\epsilon/a_{n}\right)$ can be dealt with in the same manner. Thus the proof of (ii) of Theorem 1.4 is completed.
 Remark 5.8.
 (i) Let $K(\cdot|\cdot)$ be the Kullback-Leibler divergence. Then, by computations,
-$K\left(\textup{PVII}_{m}(\theta_{1},1)|\textup{PVII}_{m}(\theta_{2},1)\right)=m(F_{m}(\nu_{1}-\nu_{2})-F_{m}(0)).$
+$K\left(\textup{PVII}_{m}(\theta_{1},1)|\textup{PVII}_{m}(\theta_{2},1)\right)=m(F_{m}(\theta_{1}-\theta_{2})-F_{m}(0)).$
 Let
-$b(\epsilon,\theta):=\inf\left\{K\left(\textup{PVII}_{m}(\theta^{\prime},1)|\textup{PVII}_{m}(\theta,1)\right)\middle||\theta^{\prime}-\theta|>\epsilon\right\}.$
+$b(\epsilon,\theta)\coloneqq\inf\left\{K\left(\textup{PVII}_{m}(\theta^{\prime},1)|\textup{PVII}_{m}(\theta,1)\right)\middle||\theta^{\prime}-\theta|>\epsilon\right\}.$
 Since $F_{m}$ is symmetric and $t\mapsto F_{m}(|t|)$ is increasing, $b(\epsilon,\theta)=m(F_{m}(\epsilon)-F_{m}(0))$. Since $F_{m}^{\prime}=-2G_{m}$,
 $\lim_{\epsilon\to+0}\frac{b(\epsilon,\theta)}{\epsilon^{2}}=\frac{m(2m-1)}{2(m+1)}=\frac{1}{I(\theta)}.$
-(ii) For the case that $m=1$, the Bahadur efficiency for the joint estimation of the location and the scale is established in [1, Theorem 4] when both the location and the scale are unknown.
+(ii) For the case that $m=1$, the Bahadur efficiency for the joint estimation of the location and the scale is established in [2, Theorem 4] when both the location and the scale are unknown.
+Recently, Akahira [1] showed that the MLE of the location parameter is first order large deviation efficient, which implies the Bahadur efficiency.
 (iii) Gao [12] obtained moderate deviation results for the maximum likelihood estimator in a more general framework under certain regular conditions.
@@
 6. Proof of Theorem 1.5
-We deal with $P\left(\hat{\theta}_{n}>r\right)$. We see that for every $r>0$ and every $\delta>0$,
-$P\left(\hat{\theta}_{n}>r\right)\leq P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)+P\left(L_{n}(0)\geq F_{m}(0)+\delta\right).$ (6.1)
-We derive upper bounds of these probabilities by modifying the assertions in Section 5. The main difference is $\delta=\delta_{r}$ diverges when $r$ tends to infinity.
-First, we give a lemma similar to Lemma 5.2. The proof differs in part. Recall that $\lambda_{m}=\frac{1}{2}(m-\frac{1}{2})$.
+We first give an outline of the proof.
+We estimate the tail probability of $\hat{\theta}_{n}$ by comparing $L_{n}(t)$ at large $|t|$ with its expectation at the true parameter $t=0$, which is $F_{m}(0)$.
+The probability of the event $\{\hat{\theta}_{n}>r\}$ is controlled by the decomposition (6.1) below.
+More specifically, if $\hat{\theta}_{n}>r$, then either $L_{n}(t)$ becomes unexpectedly small somewhere on $[r,\infty)$, or $L_{n}(0)$ becomes unexpectedly large.
+In order to bound this probability, we control the two terms in (6.1) separately by modifying the assertions in Section 5.
+Lemma 6.1 below gives a polynomial-type estimate for the lower tail of $L_{n}(t)$ at a fixed large $t$ by the exponential Chebyshev inequality.
+Then Lemma 6.2 below upgrades this pointwise estimate to the whole tail region $[r,\infty)$ and controls the first term of (6.1), by discretizing the region and using the Lipschitz bound for $L_{n}^{\prime}(t)$.
+Finally Lemma 6.3 below provides an upper-tail bound for $L_{n}(0)$ and controls the second term of (6.1).
+By symmetry, we deal with $P\left(\hat{\theta}_{n}>r\right)$. We see that for every $r>0$ and every $\delta>0$,
+(6.1) $P\left(\hat{\theta}_{n}>r\right)\leq P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)+P\left(L_{n}(0)\geq F_{m}(0)+\delta\right).$
+First, we give a lemma similar to Lemma 5.2. The proof differs in part. Recall the definition of $\lambda_{m}$ in (5.2).
 Lemma 6.1.
@@
 $P\left(\sum_{i=1}^{n}(F_{m}(t)-\log(1+(X_{i}-t)^{2}))\geq n\left(F_{m}(t)-F_{m}(0)-\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)$
-$\leq\left(\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}}\right]\right)^{n}.$
+$\leq\left(\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}}\right]\right)^{n}.$
 It holds that
-$E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}}\right]$
-$E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}},\ X_{1}\geq t/2\right]+E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}},\ X_{1}<t/2\right]$
+$E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}}\right]$
+$E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}},\ X_{1}\geq t/2\right]+E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}},\ X_{1}<t/2\right]$
 $\leq P(X_{1}\geq t/2)+\left(1+\frac{t^{2}}{4}\right)^{-\lambda_{m}}=O(t^{1-2m}),\ \ t\to\infty.$
@@
 Therefore,
-$\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}}\right]=O(t^{-4\lambda_{m}/3}),\ t\to\infty.$
+$\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}}\right]=O(t^{-4\lambda_{m}/3}),\ t\to\infty.$
 This completes the proof. ∎
-Next, we give a lemma similar to Lemma 5.3. The proof is also similar.
+Next, we give a lemma similar to Lemma 5.3. The proof is also similar. Recall the definition of $\delta_{m}(r)$ in (5.3).
 Lemma 6.2.
 There exist two positive constants $r_{m,4}$ and $c_{m,12}$ and $N_{m,1}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m,4}$ and every $n\geq N_{m,1}$,
-$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\frac{F_{m}(r)-F_{m}(0)}{4}\right)\leq c_{m,12}r^{-n\lambda_{m}/2},$
+$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{m}(r)\right)\leq c_{m,12}r^{-n\lambda_{m}/2},$
 Proof.
-For notational convenience, let $\delta_{r}:=\dfrac{F_{m}(r)-F_{m}(0)}{4}$. Since $|L_{n}^{\prime}(t)|\leq 1$,
-$\left\{\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{r}\right\}\subset\bigcup_{k\geq 1}\left\{L_{n}(k\delta_{r}+r)<F_{m}(0)+2\delta_{r}\right\}$
+Since $|L_{n}^{\prime}(t)|\leq 1$,
+$\left\{\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{m}(r)\right\}\subset\bigcup_{k\geq 1}\left\{L_{n}(k\delta_{m}(r)+r)<F_{m}(0)+2\delta_{m}(r)\right\}$
 and hence, by Lemma 6.1, for every $n\geq 2/\lambda_{m}$ and every $r\geq r_{m,4}$,
-$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{r}\right)\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{r}+r)<F_{m}(0)+2\delta_{r}\right)$
-$\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{r}+r)<F_{m}(0)+\frac{F_{m}(k\delta_{r}+r)-F_{m}(0)}{2}\right)$
-$\leq c_{m,11}\sum_{k=1}^{\infty}(k\delta_{r}+r)^{-n\lambda_{m}}\leq c_{m,11}\delta_{r}\int_{r}^{\infty}x^{-n\lambda_{m}}dx\leq c_{m,11}\delta_{r}r^{1-n\lambda_{m}}.$
-By (5.1), $\delta_{r}=O(\log r),r\to\infty$ and we have the assertion. ∎
+$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{m}(r)\right)\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{m}(r)+r)<F_{m}(0)+2\delta_{m}(r)\right)$
+$\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{m}(r)+r)<F_{m}(0)+\frac{F_{m}(k\delta_{m}(r)+r)-F_{m}(0)}{2}\right)$
+$\leq c_{m,11}\sum_{k=1}^{\infty}(k\delta_{m}(r)+r)^{-n\lambda_{m}}\leq c_{m,11}\delta_{m}(r)\int_{r}^{\infty}x^{-n\lambda_{m}}dx\leq c_{m,11}\delta_{m}(r)r^{1-n\lambda_{m}}.$
+By (5.1), $\delta_{m}(r)=O(\log r),r\to\infty$ and we have the assertion. ∎
 Finally, we give a lemma similar to Lemma 5.4.
@@
 There exist positive constants $r_{m,5}$ and $c_{m,13}$ such that for every $r>r_{m,5}$ and every $n\geq 1$,
-$P\left(L_{n}(0)\geq F_{m}(0)+\delta_{r}\right)\leq r^{-nc_{m,13}}.$
+$P\left(L_{n}(0)\geq F_{m}(0)+\delta_{m}(r)\right)\leq r^{-nc_{m,13}}.$
 Proof.
 Let $c_{m,7}$ be the constant as in the proof of Lemma 5.4. Assume that $0<\lambda\leq\lambda_{m}$. Then, by the exponential Chebyshev inequality, for every $n\geq 1$,
-$P(L_{n}(0)\geq F_{m}(0)+\delta_{r})\leq\left(\exp\left(-\lambda\delta_{r}+\frac{\lambda^{2}}{2}c_{m,7}\right)\right)^{n}.$
-By (5.1), there exists a positive constant $r_{m,5}$ such that for every $r>r_{m,5}$, $2c_{m,7}\leq\log r\leq\delta_{r}$.
-Let $\lambda_{m}^{\prime}:=\min\{1,\lambda_{m}\}$.
+$P(L_{n}(0)\geq F_{m}(0)+\delta_{m}(r))\leq\left(\exp\left(-\lambda\delta_{m}(r)+\frac{\lambda^{2}}{2}c_{m,7}\right)\right)^{n}.$
+By (5.1), there exists a positive constant $r_{m,5}$ such that for every $r>r_{m,5}$, $2c_{m,7}\leq\log r\leq\delta_{m}(r)$.
+Let $\lambda_{m}^{\prime}\coloneqq\min\{1,\lambda_{m}\}$.
 Thus, for every $r>r_{m,5}$ and every $n\geq 1$,
-$P(L_{n}(0)\geq F_{m}(0)+\delta_{r})\leq\left(\exp\left(\left(-\lambda_{m}^{\prime}+\frac{(\lambda_{m}^{\prime})^{2}}{4}\right)\delta_{r}\right)\right)^{n}\leq r^{-nc_{m,13}},$
-where we let $c_{m,13}:=\lambda_{m}^{\prime}-\frac{(\lambda_{m}^{\prime})^{2}}{4}>0$. ∎
-By applying (6.1) to $\delta=\delta_{r}$, by Lemma 6.2 and Lemma 6.3, there exist positive constants $c_{m},r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
-$P\left(\hat{\theta}_{n}>r\right)\leq r^{-c_{m}n}.$
+$P(L_{n}(0)\geq F_{m}(0)+\delta_{m}(r))\leq\left(\exp\left(\left(-\lambda_{m}^{\prime}+\frac{(\lambda_{m}^{\prime})^{2}}{4}\right)\delta_{m}(r)\right)\right)^{n}\leq r^{-nc_{m,13}},$
+where we let $c_{m,13}\coloneqq\lambda_{m}^{\prime}-\dfrac{(\lambda_{m}^{\prime})^{2}}{4}>0$. ∎
+By applying (6.1) to $\delta=\delta_{m}(r)$, by Lemma 6.2 and Lemma 6.3, there exist positive constants $r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
+$P\left(\hat{\theta}_{n}>r\right)\leq r^{-c_{m,13}n}.$
 $P\left(\hat{\theta}_{n}<-r\right)$ can be dealt with in the same manner, and we obtain Theorem 1.5.
 7. Proof of Theorem 1.6
-For $M>0$, let $\phi_{M}(x):=x^{2}\wedge M^{2}$. This is bounded and continuous on $\mathbb{R}$. By Theorem 1.2,
-$\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}\phi_{M}(x)\varphi_{m}(x)dx,$ (7.1)
-where $\varphi_{m}$ is the density function of the distribution $N\left(0,\frac{m+1}{m(2m-1)}\right)$.
+We first give an outline of the proof.
+The proof is a uniform-integrability argument based on the asymptotic normality (Theorem 1.2) and the Bahadur efficiency (two estimates used in the proof of Theorem 1.4) and the tail bound (Theorem 1.5).
+We show the lower and upper bounds separately.
+The lower bound is an easy consequence of Theorem 1.2.
+In order to obtain the upper bound, it suffices to show that the family $\left\{\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}\right\}_{n}$ is uniformly integrable, which is reduced to (7.4) below.
+The key step is to rewrite the tail contribution as an integral of tail probabilities as in (7) and then split the integral into three regimes as in (7.7).
+For the small-to-moderate region and the intermediate region, we use (5.6) and (5) used in the proof of Theorem 1.4 respectively.
+For the far tail region, we apply Theorem 1.5.
+For $M>0$, let $\phi_{M}(x)\coloneqq\min\{x^{2},M^{2}\}$.
+This is bounded and continuous on $\mathbb{R}$.
+Recall that $\varphi_{m}$ is the density function of the distribution $\displaystyle N\left(0,\dfrac{m+1}{m(2m-1)}\right)$.
+By Theorem 1.2,
+(7.1) $\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}\phi_{M}(x)\varphi_{m}(x)dx.$
 Since $x^{2}\geq\phi_{M}(x)$,
@@
 By the monotone convergence theorem,
-$\liminf_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\geq\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx.$ (7.2)
+(7.2) $\liminf_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\geq\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx=\frac{m+1}{m(2m-1)}.$
 We will show that
-$\limsup_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\leq\frac{m+1}{m(2m-1)}.$ (7.3)
+(7.3) $\limsup_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\leq\frac{m+1}{m(2m-1)}.$
 By (7.1) and the monotone convergence theorem,
-$\lim_{M\to\infty}\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx.$
-Hence, it suffices to show that
-$\limsup_{M\to\infty}\left(\limsup_{n\to\infty}E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]\right)=0.$ (7.4)
-By Fubini’s theorem for non-negative measurable functions and the change of variables $t=\sqrt{s}$,
-$E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-M^{2},\ \left|\sqrt{n}\hat{\theta}_{n}\right|\geq M\right]$
-$2\int_{M}^{\infty}tP\left(\left|\sqrt{n}\hat{\theta}_{n}\right|>t\right)dt=2n\int_{M/\sqrt{n}}^{\infty}sP\left(\left|\hat{\theta}_{n}\right|>s\right)ds.$
-We consider $P\left(\hat{\theta}_{n}>s\right)$.
-By (5.7), there exists $\epsilon_{m,3}\in(0,r_{m})$ such that for every $\epsilon\in(0,2\epsilon_{m,3})$,
-$\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\geq\frac{m(2m-1)}{4(m+1)}\epsilon^{2}.$ (7.5)
-Now we decompose the last integral into three parts:
-$\int_{M/\sqrt{n}}^{\infty}=\int_{M/\sqrt{n}}^{\epsilon_{m,3}}+\int_{\epsilon_{m,3}}^{r_{m}+1}+\int_{r_{m}+1}^{\infty},$
+$\lim_{M\to\infty}\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx=\frac{m+1}{m(2m-1)}.$
+Hence it suffices to show that
+(7.4) $\limsup_{M\to\infty}\left(\limsup_{n\to\infty}E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]\right)=0.$
+By Fubini’s theorem for non-negative measurable functions and the change of variables $t=\sqrt{s}$, we obtain that
+$\displaystyle E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]$ $\displaystyle=E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-M^{2},\ \left|\sqrt{n}\hat{\theta}_{n}\right|\geq M\right]$
+$\displaystyle=2\int_{M}^{\infty}tP\left(\left|\sqrt{n}\hat{\theta}_{n}\right|>t\right)dt$
+(7.5) $\displaystyle=2n\int_{M/\sqrt{n}}^{\infty}sP\left(\left|\hat{\theta}_{n}\right|>s\right)ds.$
+By symmetry, we consider $P\left(\hat{\theta}_{n}>s\right)$. By (5.10), there exists $\epsilon_{m,3}\in(0,r_{m})$ such that for every $\epsilon\in(0,2\epsilon_{m,3})$,
+(7.6) $\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\geq\frac{m(2m-1)}{4(m+1)}\epsilon^{2}.$
+Now we decompose the last integral in (7) into three parts:
+(7.7) $\int_{M/\sqrt{n}}^{\infty}=\int_{M/\sqrt{n}}^{\epsilon_{m,3}}+\int_{\epsilon_{m,3}}^{r_{m}+1}+\int_{r_{m}+1}^{\infty},$
 where $r_{m}$ is the constant in Theorem 1.5.
 By (7.6), (5), and (5.6), there exist two positive constants $c_{m,14},c_{m,15}$ and $N_{m,2}\in\mathbb{N}$ depending only on $m$ such that for every $n\geq N_{m,2}$ and $s\in(0,2\epsilon_{m,3})$,
-$P\left(\hat{\theta}_{n}>s\right)\leq\exp\left(-\frac{m(2m-1)}{4(m+1)}s^{2}n\right)+c_{m,14}\exp(-c_{m,15}n).$
+(7.8) $P\left(\hat{\theta}_{n}>s\right)\leq\exp\left(-\frac{m(2m-1)}{4(m+1)}s^{2}n\right)+c_{m,14}\exp(-c_{m,15}n).$
 Therefore, for $n\geq N_{m,2}$,
@@
 Hence,
-$\limsup_{n\to\infty}2n\int_{M/\sqrt{n}}^{\epsilon_{m,3}}sP\left(\hat{\theta}_{n}>s\right)ds\leq\frac{4(m+1)}{m(2m-1)}\exp\left(-\frac{m(2m-1)}{4(m+1)}M^{2}\right).$ (7.6)
+(7.9) $\limsup_{n\to\infty}2n\int_{M/\sqrt{n}}^{\epsilon_{m,3}}sP\left(\hat{\theta}_{n}>s\right)ds\leq\frac{4(m+1)}{m(2m-1)}\exp\left(-\frac{m(2m-1)}{4(m+1)}M^{2}\right).$
 Since
 $2n\int_{\epsilon_{m,3}}^{r_{m}+1}sP\left(\hat{\theta}_{n}>s\right)ds\leq 2(r_{m}+1)^{2}nP\left(\hat{\theta}_{n}>\epsilon_{m,3}\right),$
-$\limsup_{n\to\infty}2n\int_{\epsilon_{m,3}}^{r_{m}+1}sP\left(\hat{\theta}_{n}>s\right)ds=0.$ (7.7)
+by applying (7.8) to $s=\epsilon_{m,3}$,
+(7.10) $\limsup_{n\to\infty}2n\int_{\epsilon_{m,3}}^{r_{m}+1}sP\left(\hat{\theta}_{n}>s\right)ds=0.$
 By Theorem 1.5, for large $n$,
@@
 Hence,
-$\limsup_{n\to\infty}2n\int_{r_{m}+1}^{\infty}sP\left(\hat{\theta}_{n}>s\right)ds=0.$ (7.8)
+(7.11) $\limsup_{n\to\infty}2n\int_{r_{m}+1}^{\infty}sP\left(\hat{\theta}_{n}>s\right)ds=0.$
 By (7.9), (7.10), and (7.11),
@@
 The same estimate holds for $P\left(\hat{\theta}_{n}<-s\right)$. Since the right hand side converges to $0$ as $M\to\infty$, (7.4) holds.
-Thus we obtain (7.2) and (7.3) and the proof is completed.
+Thus we obtain (7.2) and (7.3), and the proof of Theorem 1.6 is completed. We provide numerical evidence for Theorem 1.6 in Subsection 8.4 below.
 Remark 7.1.
-The variance of the maximum likelihood estimator of the parameter $m$ was dealt with by Taylor’s unpublished manuscript [14].
-8. Numerical computations
-We perform simulation studies using the software R to illustrate the properties of the maximum likelihood estimator.
-We used R version 4.5.1.
-We deal with $nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$ appearing in Theorem 1.6.
-We can assume that $\theta=0$ without loss of generality.
-We consider parameters $m=0.1\times k$ for $6\leq k\leq 15$ and sample sizes $n=10,50,100,500,1000$.
-In each choice of the pair $(m,n)$, we compute the average of $n\left(\hat{\theta}_{n}\right)^{2}$ over $10^{7}$ samples of size $n$ generated by the rpearsonVII() function in the package ‘PearsonDS’, and in the optimization, we use the nlminb() function with the starting point being the median.
-Table 1 gives the result.
-In the case of $m=1$, [6, Table 2] gives numerical computations for $n=5,6,\dots,14,15,20,25,\dots,35,40$.
-This is consistent with the numerical computation in [1, Table 2] for the joint estimation of the location and scale.
-The convergence becomes faster as the parameter $m$ increases.
-Since $\hat{\theta}_{n}$ appears not to be square-integrable if $m=0.6$ and $n=10$, and the numerical computation is not stable although $\hat{\theta}_{n}$ appears to be square-integrable if $m=0.7$ and $n=10$, we mark them as not available (NA).
-By this table, we conjecture that for each $m$, $\left(nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]\right)_{n}$ is decreasing in $n$.
-[…]
-0.6 NA 25.756 16.108 13.756 13.545 13.333
-0.7 NA 7.236 6.566 6.160 6.116 6.071
-0.8 9.046 4.156 3.935 3.783 3.766 3.750
-0.9 4.524 2.832 2.729 2.655 2.648 2.639
-1 2.908 2.108 2.052 2.011 2.006 2.000
-1.1 2.103 1.659 1.624 1.599 1.595 1.591
-1.2 1.630 1.356 1.332 1.314 1.311 1.310
-1.3 1.320 1.138 1.123 1.108 1.107 1.106
-1.4 1.105 0.977 0.964 0.954 0.953 0.952
-1.5 0.945 0.851 0.842 0.835 0.834 0.833
-Table 1.
-Simulated values of $n\,\mathbb{E}\!\left[(\hat{\theta}_{n}-\theta)^{2}\right]$ ($\theta=0$).
-Rows correspond to $m$; columns to the sample size $n$.
-The column labeled $\infty$ reports the theoretical limit (m+1)/(m(2m1))(m+1)/(m(2m-1)) given by Theorem 1.6.
-We round the results to three decimal places.
+(i) The variance of the maximum likelihood estimator of the parameter $m$ was dealt with by Taylor’s unpublished manuscript [20].
+[14, pp.396–399] discusses the parameter estimation other than the location.
+(ii) By symmetry, we strongly expect that $\hat{\theta}_{n}$ is unbiased, that is, $E\left[\hat{\theta}_{n}\right]=\theta$ for each $n$ such that $\hat{\theta}_{n}$ is integrable.
+However, to the best of our knowledge, no rigorous proof of this fact has been established.
+We provide numerical evidence for this in Subsection 8.1 below.
+8. Simulation study
+We perform simulation studies using R to illustrate the properties of the maximum likelihood estimator.
+We consider parameter values $m=0.1\cdot k$ for $6\leq k\leq 15$ and sample sizes $n=10,50,100,500,1000$.
+In the tables below, the rows correspond to $m$ and the columns to the sample size $n$, unless otherwise stated.
+In the simulations, we let $\theta=0$.
+We use R version 4.5.2.
+For each pair $(m,n)$, we generate $N=10^{7}$ samples of size $n$ using the rpearsonVII() function in the package ‘PearsonDS’, and compute the corresponding MLEs.
+For the optimization, we use the nlminb() function with the sample median as the starting value.
+8.1. Bias
+We compute $\left|E\left[\hat{\theta}_{n}-\theta\right]\right|$.
+We approximate $E\left[\hat{\theta}_{n}-\theta\right]$ by $\displaystyle\frac{1}{N}\sum_{i=1}^{N}Z_{i}$, where $(Z_{i})_{i}$ are i.i.d.
+random variables distributed as $\hat{\theta}_{n}-\theta$.
+$\n\displaystyle m\backslash$ n 10 50 100 500 1000
+0.6 2.3 $3.7\cdot 10^{-5}$ $8.5\cdot 10^{-5}$ $1.1\cdot 10^{-4}$ $4.4\cdot 10^{-5}$
+0.7 $7.8\cdot 10^{-4}$ $1.7\cdot 10^{-5}$ $1.2\cdot 10^{-4}$ $2.6\cdot 10^{-5}$ $2.9\cdot 10^{-6}$
+0.8 $3.3\cdot 10^{-5}$ $2.9\cdot 10^{-5}$ $1.9\cdot 10^{-5}$ $3.2\cdot 10^{-5}$ $5.6\cdot 10^{-6}$
+0.9 $7.3\cdot 10^{-5}$ $7.1\cdot 10^{-5}$ $5.2\cdot 10^{-5}$ $7.2\cdot 10^{-6}$ $3.9\cdot 10^{-5}$
+1.0 $1.8\cdot 10^{-4}$ $5.2\cdot 10^{-5}$ $6.2\cdot 10^{-5}$ $4.5\cdot 10^{-6}$ $2.3\cdot 10^{-5}$
+1.1 $8.6\cdot 10^{-6}$ $6.5\cdot 10^{-5}$ $2.0\cdot 10^{-5}$ $5.2\cdot 10^{-6}$ $1.2\cdot 10^{-5}$
+1.2 $6.0\cdot 10^{-5}$ $6.9\cdot 10^{-5}$ $4.3\cdot 10^{-5}$ $1.4\cdot 10^{-5}$ $7.6\cdot 10^{-6}$
+1.3 $3.7\cdot 10^{-6}$ $3.0\cdot 10^{-6}$ $8.9\cdot 10^{-5}$ $4.5\cdot 10^{-6}$ $3.2\cdot 10^{-6}$
+1.4 $5.7\cdot 10^{-5}$ $1.8\cdot 10^{-5}$ $2.7\cdot 10^{-5}$ $6.3\cdot 10^{-6}$ $4.7\cdot 10^{-6}$
+1.5 $4.2\cdot 10^{-5}$ $2.8\cdot 10^{-5}$ $6.3\cdot 10^{-6}$ $1.6\cdot 10^{-5}$ $1.5\cdot 10^{-6}$
+Table 1. Simulated values of $\left|E\left[\hat{\theta}_{n}-\theta\right]\right|$.
+This table suggests that $\hat{\theta}_{n}$ is not integrable for $m=0.6$ and $n=10$.
+In this case, it is interesting to compute $P\left(\hat{\theta}_{n}-\theta>r\right)$ for large $r$.
+We approximate this probability by the proportion of simulated values which are larger than $r$ among the $N=10^{7}$ simulated observations.
+$r$ $10^{3}$ $5\cdot 10^{3}$ $10^{4}$ $5\cdot 10^{4}$ $10^{5}$ $5\cdot 10^{5}$ $10^{6}$
+$-1.22$ $-1.08$ $-1.02$ $-0.94$ $-0.94$ $-0.95$ $-0.96$
+Table 2. Simulated values of $\dfrac{\log P(\hat{\theta}_{n}-\theta>r)}{\log r}$.
+This table suggests that there exists $C>0$ such that $P\left(\hat{\theta}_{n}-\theta>r\right)\geq Cr^{-1}$ for large $r$, which implies $\hat{\theta}_{n}$ is not integrable.
+8.2. Asymptotic normality
+Recall that $\varphi_{m}$ is the density of the normal distribution $N\left(0,\dfrac{m+1}{m(2m-1)}\right)$. We consider the Kolmogorov–Smirnov metric:
+$\Delta_{n}\coloneqq\sup_{x\in\mathbb{R}}\left|P\left(\hat{\theta}_{n}-\theta\leq x\right)-\int_{-\infty}^{x}\varphi_{m}(t)dt\right|.$
+We approximate $P\left(\hat{\theta}_{n}-\theta\leq x\right)$ by the empirical cumulative distribution function $\displaystyle\frac{1}{N}\sum_{i=1}^{N}{\bf 1}_{\{Z_{i}\leq x\}}$ where $(Z_{i})_{i}$ are i.i.d.
+random variables distributed as $\hat{\theta}_{n}-\theta$.
+We recall the Dvoretzky-Kiefer-Wolfowitz-Massart inequality [9, Theorem 1.32]:
+$P\left(\sup_{x\in\mathbb{R}}\left|\frac{1}{N}\sum_{i=1}^{N}{\bf 1}_{\{Z_{i}\leq x\}}-P\left(\hat{\theta}_{n}-\theta\leq x\right)\right|>\epsilon\right)\leq 2\exp(-2N\epsilon^{2}).$
+$\n\displaystyle m\backslash$ n 10 50 100 500 1000
+0.6 $1.1\cdot 10^{-1}$ $2.4\cdot 10^{-2}$ $1.2\cdot 10^{-2}$ $2.4\cdot 10^{-3}$ $1.3\cdot 10^{-3}$
+0.7 $5.3\cdot 10^{-2}$ $1.1\cdot 10^{-2}$ $5.5\cdot 10^{-3}$ $1.2\cdot 10^{-3}$ $5.8\cdot 10^{-4}$
+0.8 $3.4\cdot 10^{-2}$ $6.9\cdot 10^{-3}$ $3.5\cdot 10^{-3}$ $7.9\cdot 10^{-4}$ $4.7\cdot 10^{-4}$
+0.9 $2.5\cdot 10^{-2}$ $5.0\cdot 10^{-3}$ $2.5\cdot 10^{-3}$ $6.4\cdot 10^{-4}$ $3.5\cdot 10^{-4}$
+1.0 $1.9\cdot 10^{-2}$ $4.0\cdot 10^{-3}$ $2.0\cdot 10^{-3}$ $6.4\cdot 10^{-4}$ $2.9\cdot 10^{-4}$
+1.1 $1.5\cdot 10^{-2}$ $3.1\cdot 10^{-3}$ $1.5\cdot 10^{-3}$ $4.2\cdot 10^{-4}$ $3.1\cdot 10^{-4}$
+1.2 $1.3\cdot 10^{-2}$ $2.6\cdot 10^{-3}$ $1.4\cdot 10^{-3}$ $3.0\cdot 10^{-4}$ $2.7\cdot 10^{-4}$
+1.3 $1.1\cdot 10^{-2}$ $2.3\cdot 10^{-3}$ $1.2\cdot 10^{-3}$ $3.4\cdot 10^{-4}$ $3.2\cdot 10^{-4}$
+1.4 $9.3\cdot 10^{-3}$ $1.9\cdot 10^{-3}$ $9.7\cdot 10^{-4}$ $2.8\cdot 10^{-4}$ $2.1\cdot 10^{-4}$
+1.5 $8.2\cdot 10^{-3}$ $1.6\cdot 10^{-3}$ $9.3\cdot 10^{-4}$ $3.8\cdot 10^{-4}$ $2.4\cdot 10^{-4}$
+Table 3. Simulated values of $\Delta_{n}$.
+8.3. Confidence intervals
+We consider the pivotal quantity $\sqrt{n}\left(\hat{\theta}_{n}-\theta\right)$ of the model.
+Let $z_{\beta}$ denote the upper $\beta$-quantile, that is, the value satisfying $P\left(\sqrt{n}(\hat{\theta}_{n}-\theta)\geq z_{\beta}\right)=\beta$ for $0<\beta<\dfrac{1}{2}$.
+We report the values of $z_{\alpha/2}$ for $\alpha=0.1,0.05,0.01$.
+We approximate $z_{\alpha/2}$ by sorting $N=10^{7}$ MLEs and using the quantile function in R.
+Using $z_{\alpha/2}$, we obtain the $100(1-\alpha)\%$ confidence interval for the location parameter $\theta$:
+$\left[\hat{\theta}_{n}-\frac{z_{\alpha/2}}{\sqrt{n}},\hat{\theta}_{n}+\frac{z_{\alpha/2}}{\sqrt{n}}\right].$
+In the following tables, the column labeled $\infty$ reports the upper $\alpha/2$ quantile of the limiting normal distribution $N\left(0,\dfrac{m+1}{m(2m-1)}\right)$ given by Theorem 1.2.
+$\n\displaystyle m\backslash$ n 10 50 100 500 1000 $\infty$
+0.6 19.60 7.06 6.46 6.08 6.05 6.01
+0.7 6.34 4.34 4.19 4.08 4.06 4.05
+0.8 4.09 3.33 3.25 3.20 3.19 3.19
+0.9 3.17 2.75 2.71 2.68 2.68 2.67
+1.0 2.64 2.38 2.35 2.33 2.33 2.33
+1.1 2.29 2.11 2.10 2.08 2.08 2.07
+1.2 2.04 1.91 1.90 1.89 1.88 1.88
+1.3 1.85 1.75 1.74 1.73 1.73 1.73
+1.4 1.70 1.62 1.61 1.61 1.61 1.61
+1.5 1.58 1.52 1.51 1.50 1.50 1.50
+Table 4. Simulated values of $z_{\alpha/2}$ ($\alpha=0.1$).
+$\n\displaystyle m\backslash$ n 10 50 100 500 1000 $\infty$
+0.6 37.96 9.14 7.96 7.29 7.22 7.16
+0.7 9.46 5.34 5.07 4.87 4.85 4.83
+0.8 5.57 4.04 3.91 3.82 3.80 3.80
+0.9 4.12 3.33 3.25 3.20 3.19 3.18
+1.0 3.35 2.87 2.82 2.78 2.78 2.77
+1.1 2.87 2.54 2.51 2.48 2.48 2.47
+1.2 2.53 2.29 2.27 2.25 2.25 2.24
+1.3 2.28 2.10 2.08 2.06 2.06 2.06
+1.4 2.09 1.94 1.93 1.92 1.92 1.91
+1.5 1.93 1.82 1.80 1.79 1.79 1.79
+Table 5. Simulated values of $z_{\alpha/2}$ ($\alpha=0.05$).
+$\n\displaystyle m\backslash$ n 10 50 100 500 1000 $\infty$
+0.6 162.34 15.48 11.47 9.73 9.57 9.41
+0.7 21.55 7.66 6.91 6.45 6.39 6.35
+0.8 10.48 5.58 5.27 5.04 5.01 4.99
+0.9 6.99 4.53 4.35 4.21 4.20 4.18
+1.0 5.31 3.87 3.75 3.67 3.65 3.64
+1.1 4.35 3.41 3.33 3.27 3.26 3.25
+1.2 3.74 3.07 3.01 2.96 2.95 2.95
+1.3 3.29 2.80 2.75 2.72 2.71 2.71
+1.4 2.97 2.59 2.55 2.52 2.52 2.51
+1.5 2.71 2.41 2.38 2.36 2.35 2.35
+Table 6. Simulated values of $z_{\alpha/2}$ ($\alpha=0.01$).
+In the case $m=1$, [11, Section 4] provides a table of quantiles $z_{\alpha/2}$ for sample sizes $n=5,10,\dots,40$ and $\alpha=0.1,0.05,0.01$.
+8.4. Asymptotic efficiency
+We consider the quantity $nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$ appearing in Theorem 1.6.
+The column labeled $\infty$ reports the theoretical limit $\dfrac{m+1}{m(2m-1)}$ given by Theorem 1.6.
+$\n\displaystyle m\backslash$ n 10 50 100 500 1000 $\infty$
+0.6 NA 25.756 16.108 13.756 13.537 13.333
+0.7 NA 7.231 6.561 6.162 6.114 6.071
+0.8 8.979 4.154 3.933 3.786 3.769 3.750
+0.9 4.511 2.831 2.730 2.657 2.648 2.639
+1.0 2.907 2.110 2.053 2.011 2.005 2.000
+1.1 2.106 1.660 1.623 1.598 1.595 1.591
+1.2 1.631 1.356 1.332 1.314 1.311 1.310
+1.3 1.322 1.138 1.122 1.108 1.107 1.106
+1.4 1.105 0.976 0.964 0.955 0.953 0.952
+1.5 0.946 0.851 0.842 0.835 0.834 0.833
+Table 7. Simulated values of $n\,\mathbb{E}\!\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$.
+In the case $m=1$, [11, Table 2] provides numerical values for $n=5,6,\dots,14,15,20,25,\dots,35,40$.
+These results are consistent with the numerical results in [2, Table 2] for the joint estimation of the location and scale.
+Acknowledgements. The author would like to express his gratitude to the referees for their helpful comments and suggestions.